### v6

# Establishing a Foundation for Tetun Ad-hoc Text Retrieval: Stemming, Indexing, Retrieval, and Ranking

Abstract.
Searching for information on the internet and digital platforms requires effective retrieval solutions. However, such solutions are not yet available for Tetun, making it difficult to find relevant documents for search queries in this language. To address this gap, we investigate Tetun text retrieval with a focus on the ad-hoc retrieval task. The study begins with the development of essential language resources—including a list of stopwords, a stemmer, and a test collection—that serve as a foundation for Tetun text retrieval. Various strategies are evaluated using document titles and content. The results show that retrieving document titles, after removing hyphens and apostrophes but without applying stemming, improves performance compared to the baseline. Efficiency increases by 31.37%, while effectiveness achieves an average relative gains of +9.40% in MAP@10 and +30.35% in NDCG@10 with DFR BM25. Beyond the top-10 cutoff point, Hiemstra LM demonstrates strong performance across multiple retrieval strategies and evaluation metrics. The contributions of this work include the development of Labadain-Stopwords (a list of 160 Tetun stopwords), Labadain-Stemmer (a Tetun stemmer with three variants), and Labadain-Avaliadór (a Tetun test collection comprising 59 topics, 33,550 documents, and 5,900 qrels). These resources are publicly available to support future research in Tetun information retrieval.
Stopwords, Stemming, Test Collection, Ad-hoc Retrieval
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information Retrieval Text Retrievalccs: Low-Resource Languages Tetun
1. Introduction
Ad-hoc text retrieval is the task of retrieving documents from large text collections in response to user queries without prior knowledge of the topics that users are likely to search, highlighting the unpredictable nature and short duration of each search (Voorhees and Harman, 1998; Voorhees, 2004). Users typically express their information needs through natural language text queries and submit them to a search system. The retrieval system then retrieves, ranks, and returns documents relevant to the query, presenting the most relevant documents at the top of the list, with less relevant ones further down.
Effective information retrieval (IR) systems are essential for accessing the extensive digital content available on the web and digital platforms. Evaluating the effectiveness of these IR systems relies on robust test collections. High-resource languages benefit from readily available test collections sourced from various publicly accessible repositories, such as the IR dataset catalog (MacAvaney et al., 2021) [^1] https://ir-datasets.com and HuggingFace. [^2] https://huggingface.co/datasets However, this scenario differs for low-resource languages (LRLs), where data scarcity and linguistic complexities make accessing test collections challenging.
The classical approach for constructing test collections follows the Cranfield paradigm (Cleverdon, 1967), which became widely recognized through the Text REtrieval Conference (TREC) series of large-scale evaluation campaigns (Harman, 1992). In this TREC-style adaptation of the Cranfield approach, a test collection comprises three components: a document collection, a set of information needs (or topics), and corresponding relevance judgments. In ad-hoc text retrieval, a set of topics is formulated and then tested by searching large document collections to estimate the number of relevant documents returned for each topic (Sanderson, 2010). These query-document pairs are then provided to assessors for relevance judgment. Traditionally, relevance judgments are made by human assessors, involving a process that is both time-intensive and costly. Due to financial constraints, relevance assessment tasks for constructing test collections for LRLs are often carried out by volunteer native language speakers, such as students (Sahu and Pal, 2023; AleAhmad et al., 2009).
To identify effective retrieval strategies, several approaches are explored and tested using a reliable test collection. The classical approach to configuring these strategies involves preprocessing documents and queries, primarily focusing on stopword removal and stemming. For stopwords removal, a readily available list of stopwords is necessary, and a proper stemmer is required to process the input text. However, these resources are often unavailable for most LRLs.
These challenges are also faced in the development of resources for Tetun, a LRL spoken by over 923,000 people in Timor-Leste (de Jesus, 2023). Timor-Leste is a Southeast Asian island country characterized by its multilingualism, comprising two official languages (Tetun and Portuguese), two working languages (English and Indonesian) (Vasconcelos et al., 2011), and over 30 dialects spoken across the territory (de Jesus, 2023). Tetun, which was a dialect, became one of Timor-Leste’s official languages when the country restored its independence in 2002 (Vasconcelos et al., 2011). Despite its status as an official language, Tetun is characterized by data scarcity, with fewer than 45,000 documents available on the web as of 2023 (Kudugunta et al., 2023; de Jesus and Nunes, 2024a). Moreover, Tetun is a less-studied and computerized language, lacking essential resources for effective text retrieval, including a stopword list, a stemmer, and a test collection for the ad-hoc retrieval task.
To tackle the aforementioned challenges, we investigated strategies for Tetun ad-hoc text retrieval, including evaluating the impact of stemming and stopwords, to identify the most effective retrieval solutions for Tetun. The research questions (RQs) we addressed in this study are the following:
RQ1
How can text preprocessing techniques tailored to Tetun’s linguistic characteristics improve retrieval effectiveness?
RQ2
What strategies provide the most effective solutions for Tetun text-based search?
Given that Tetun words contain accented letters (á, é, í, ó, ú, ñ), apostrophes (‘), and hyphens in monosemantic compound words, our objective is to investigate the impact of query and document preprocessing on the effectiveness of text retrieval in Tetun text-based search. In line with the research questions above, we hypothesize that applying language-specific preprocessing to queries and documents can improve retrieval effectiveness without the need for stemming, particularly when retrieving short texts such as document titles. This hypothesis is grounded in findings from our preliminary study on Tetun ad-hoc text retrieval, which reported a 3.1% relative improvement in overall MAP when stemming was not applied (de Jesus, 2022).
To test this hypothesis, we began by developing a list of Tetun stopwords (Labadain-Stopwords), a Tetun stemmer (Labadain-Stemmer), and a Tetun test collection (Labadain-Avaliadór) using the Labadain-30k+ dataset (de Jesus and Nunes, 2024a). For Labadain-Stemmer, three variants were developed: light, moderate, and heavy. The Labadain-Stemmer performance was evaluated both as standalone systems and for their impact within the retrieval system (intrinsic) and extrinsic assessments (Mollá and Hutchinson, 2003; Jones and Galliers, 1996). The Labadain-Avaliadór was developed following TREC guidelines and assessed by native Tetun-speaking students. This collection was then used to evaluate the retrieval effectiveness for Tetun ad-hoc text retrieval. The contributions of this work include: (i) the development of Labadain-Stopwords, (ii) the creation of Labadain-Stemmer with three variants, (iii) the construction of Labadain-Avaliadór, and (iv) the establishment of baselines for Tetun ad-hoc text retrieval.
The remainder of this paper is organized as follows. Section 2 reviews background and related work. Section 3 provides an overview of Tetun and its linguistic characteristics. The dataset used in this study is described in Section 4, and Section 5 outlines the methodology for establishing baselines in Tetun ad-hoc text retrieval. Section 6 presents the construction of Labadain-Stopwords, while Section 7 describes the development of Labadain-Stemmer. Section 8 details the creation of Labadain-Avaliadór, and Section 9 reports the baselines for Tetun ad-hoc text retrieval. Finally, Section 10 concludes the paper and discusses directions for future work.
2. Background and Related Work
Text retrieval typically involves several preprocessing steps, such as stopword removal and stemming. The key topics relevant to this study are the development of stopword lists, stemming approaches, test collections for evaluation, and baselines for Tetun ad-hoc text retrieval. Background on each of these topics is provided in the following subsections.
2.1. Stopwords
Stopwords are function words—such as articles, prepositions, and conjunctions—that appear frequently in documents. Traditionally, stopword lists are created by selecting the most frequent terms in a corpus, often choosing the top-$n$ most common terms (Croft et al., 2009; Manning et al., 2009). This approach typically relies on classical term weighting techniques such as term frequency (TF) (Luhn, 1957), inverse document frequency (IDF) (Sparck Jones, 1972), or term frequency-inverse document frequency (TF-IDF) (Salton et al., 1975).
Fox (1990) applied a term frequency approach to the Brown Corpus to generate a stopword list for English, a method that has since been widely adopted to develop stopwords for other languages, including French (Savoy, 1999); Marathi, Czech, Hungarian, and five other languages (Ghosh and Bhattacharya, 2017); and Kinyarwanda and Kirundi (Niyongabo et al., 2020).
Furthermore, Lo et al. (2005) introduced the normalized inverse document frequency (NIDF) for stopword detection, evaluating it on four English TREC collections and showing that NIDF outperforms TF and IDF. Later, Ferilli (Ferilli, 2021) proposed the term-document frequency (TDF) metric and, after testing it on two Italian corpora, found that TDF surpasses TF, IDF, and NIDF, particularly in smaller datasets. More recently, Ali et al. (2024) demonstrated that a network-based approach exploiting topological properties of co-occurrence networks—such as in-degree, out-degree, and degree—outperforms traditional term-weighting techniques, with in-degree yielding the most consistent results across both high- and low-resource languages, including Tetun.
In IR, stopwords generally contribute minimal value to retrieving relevant documents for a given query (Baeza-Yates and Ribeiro-Neto, 2011; Manning et al., 2009). Therefore, removing these words from both queries and documents can enhance retrieval efficiency and effectiveness (Sahu et al., 2023; Sahu and Pal, 2023; Baeza-Yates and Ribeiro-Neto, 2011; Croft et al., 2009). Despite this, other studies have reported that the effectiveness of stopword removal varies between ranking models and languages, demonstrating that it is beneficial in some cases but not in others (Ghosh and Bhattacharya, 2017; Dolamic and Savoy, 2010; Savoy, 1999). In a study on French, Savoy (1999) found that retaining stopwords performed better than removing them when using BM25. Similarly, Dolamic and Savoy (2010) did not observe significant differences in retrieval effectiveness for Marathi and Bengali, while for Hindi, including stopwords improved retrieval effectiveness, with average gains of approximately 20% in mean average precision (MAP). Ghosh and Bhattacharya (2017) further demonstrated variability in retrieval effectiveness across datasets within the same language, observing that stopword removal did not lead to noticeable differences in retrieval effectiveness when evaluated on datasets from FIRE, CLEF, and TREC collections across several languages.
2.2. Stemming
Stemming is an essential component of text processing that captures the relationship between different variations and forms of a word resulting from inflection (e.g., plurals, tenses, and gender) or derivation (e.g., converting a verb into a noun by adding suffixes) and reduces them to a common root (Croft et al., 2009). A stem is the root form of a word that remains after the removal of its affixes (Baeza-Yates and Ribeiro-Neto, 2011). Stemming can be useful for improving retrieval effectiveness by minimizing index size and reducing the number of distinct terms.
The first stemming algorithm was proposed by Lovins (1968), based on the principles of iteration and longest match. The iteration principle assumes that affixes are attached to stems in a particular order from a predefined set of affixes. The algorithm removes affixes from either the beginning or the end of the word, depending on which affix class is detected. According to the longest match principle, if multiple endings within a class match, the longest one should be removed. Since then, various stemming techniques have been developed, including rule-based, dictionary-based, and automatic stemmers. One of the most notable examples is the Porter Stemmer (Porter, 1980), a widely used rule-based suffix removal algorithm for English stemming due to its simplicity and performance (Baeza-Yates and Ribeiro-Neto, 2011).
Several Asian languages have adopted suffix-stripping stemmers based on the Porter and Lovins approaches, including Sanskrit (Sahu and Pal, 2023), Sundanese (Ardiyanti et al., 2018), Czech (Dolamic and Savoy, 2009), and Indonesian (Adriani et al., 2007). An advanced version of the Porter Stemmer is the Snowball stemming algorithm, which supports multiple languages, including Portuguese, Spanish, and German (Snowball, 2002). Snowball applies a set of predefined stemming rules tailored to the specific morphological structure of each language, primarily focusing on removing suffixes.
In an experiment on French, Savoy (1999) found that stemming was particularly beneficial for retrieving short documents (scientific abstracts averaging 24.5 indexing terms), with improvements in average precision (AP) in various retrieval models. For longer documents (news articles averaging 182.2 terms per article) or cases where accents were ignored, stemming yielded only marginal benefits. Similarly, Braschler and Ripplinger (2004), evaluating on German data from the CLEF 2000 and 2001 datasets, reported stemming gains in MAP of up to 23% for short queries (title only) and up to 11% for long queries (a combination of title, description, and narrative).
However, Hollink et al. (2004), in a monolingual document retrieval experiment using the CLEF 2002 dataset across eight languages, reported inconsistent results. Specifically, stemming improved MAP for Finnish, French, German, and Swedish but had no positive effect on Dutch, English, Italian, or Spanish. Likewise, Flores and Moreira (2016), in an experiment with four different languages from the CLEF 2005 and 2006 datasets, reported that stemming was generally beneficial in MAP for Portuguese, French, and Spanish but not for English when tested with different stemmers.
In studies on Asian LRLs, Sahu and Pal (2023) reported that stemming improved retrieval effectiveness for Sanskrit by 4.31% in MAP across multiple ranking models. Likewise, Sahu et al. (2023) observed a 1.41% MAP improvement for Urdu when testing several stemming approaches, while Adriani et al. (2007) found a 2.00% MAP improvement in experiments with Indonesian.
2.3. Test Collection
A reliable test collection is essential for evaluating the effectiveness of retrieval systems. For high-resource languages, these collections are typically made available through large-scale campaigns such as the TREC, [^3] https://trec.nist.gov the Conference and Labs of the Evaluation Forum (CLEF), [^4] https://www.clef-initiative.eu the NII Testbeds and Community for Information Access Research project (NTCIR), [^5] http://research.nii.ac.jp/ntcir/index-en.html and the Forum for Information Retrieval Evaluation (FIRE). [^6] http://fire.irsi.res.in/ Following the Cranfield paradigm implemented in the TREC, developing a test collection involves selecting various retrieval strategies to compare and produce top-ranked lists of documents (runs). These runs are then merged to create a pooled set of documents for each query. This pool is manually judged for relevance by human assessors, producing a list of relevant documents (qrels(Sanderson, 2010).
Relevance judgments typically fall into two categories: binary and graded relevance. Binary relevance categorizes each document as either relevant or non-relevant to the user’s query, assigning a score of 1 for relevant and 0 for non-relevant documents. The graded relevance evaluates documents on multiple levels of relevance, with the most relevant documents awarding higher scores. Binary relevance is predominantly used for experimental research in the TREC collections. In the TREC-9 Web Track, three-level graded relevance was introduced: not relevant, relevant, and highly relevant (Hawking, 2000). Later, Sormunen proposed a four-level relevance scale consisting of non-relevant, marginally relevant, relevant, and highly relevant (Sormunen, 2002). Kekäläinen (2005) adopted a similar four-point scale but labeled the third level as “fairly relevant” rather than “relevant”. This four-point scale was subsequently implemented across multiple TREC tracks. The ad-hoc retrieval task was a central focus of the TREC tracks held from 1992 to 1999 and was revisited on the robust track from 2003 to 2005 (Voorhees, 2006).
The TREC-style approach, derived from the Cranfield paradigm, is commonly used to develop test collections for LRLs. Sahu and Pal (2023) applied this method to create a Sanskrit test collection comprising 7,057 news articles and 50 topics, with queries and relevance judgments produced by two Ph.D. students. Similarly, Chavula and Suleman (2021) constructed a test collection for three Bantu languages—Chichewa, Citumbuka, and Cinyanja—using documents from newspapers, Wikipedia, and web pages. Their collection includes 13,627 documents and 387 topics, with queries and relevance assessments carried out by six recruited assessors. Furthermore, AleAhmad et al. (2009) developed the Hamshahri test collection for Persian, based on a news corpus of 166,774 documents and 65 queries, with queries and relevance assessments performed by 17 volunteer students.
2.4. Summary
In ad-hoc text retrieval, preprocessing steps, such as stopword removal and stemming, are often employed to enhance retrieval efficiency and effectiveness. However, studies show that the impact of these techniques on retrieval effectiveness varies across languages, proving beneficial in some instances but less so in others. Furthermore, Ghosh and Bhattacharya (2017) highlighted that the influence of stopwords can differ even within the same language across different collections, such as Bangla and Hindi in the FIRE datasets of 2010 and 2011.
The evaluation of retrieval system effectiveness relies on robust test collections, which are typically developed following TREC guidelines, with human assessors conducting relevance judgments. For LRLs, the same methodologies are adapted to create test collections. However, due to financial constraints, relevance assessments in these less-resourced contexts are often carried out by students who are native language speakers.
This study addresses a critical gap in Tetun text retrieval by introducing three essential resources: a stopword list (Labadain-Stopwords), a language-specific stemmer (Labadain-Stemmer), and a Tetun test collection (Labadain-Avaliadór). Through a series of experiments, we evaluate various retrieval strategies to establish baselines and identify the most effective approach for ad-hoc text retrieval in Tetun. The subsequent sections provide a detailed overview of the development of each resource and its application in the experiments.
3. Tetun
This section presents an overview of Tetun, including its orthography, morphology, and Portuguese loanwords.
3.1. Overview
Tetun, alternatively written as Tetum (in English) or Tétum (in Portuguese), is an Austronesian language spoken in Timor-Leste, an island nation in Southeast Asia. It has two primary varieties: Tetun Dili, also known as Tetun Prasa (commonly referred to simply as Tetun), and Tetun Terik (van Klinken et al., 2002). Tetun has two standardized forms: one developed by the Instituto Nacional de Linguística (Tetun INL) and another by the Dili Institute of Technology (Tetun DIT). Tetun Terik, meanwhile, remains one of the dialects spoken in Timor-Leste.
Tetun is one of Timor-Leste’s official languages alongside Portuguese (Vasconcelos et al., 2011). The government recognized Tetun INL as the official Tetun, which is used in the education system, official publications, and media (DL 01/2004, 2004). Tetun DIT was developed by linguists at the Dili Institute of Technology with some standardized differences from Tetun INL in terms of writing conventions (van Klinken et al., 2002). For example, the words [  (give), ne’ebé (which/that) ] in Tetun INL correspond to [ foo, neebe ] in Tetun DIT. According to the 2015 census report, Timor-Leste’s population was 1.18 million, with 78.78% of the population being Tetun speakers (de Jesus and Nunes, 2024a). Among them, 30.50% considered Tetun their home language, while 48.28% spoke it as a second or third language. Census 2022 reported a population growth of 13.40%, increasing from 1.18 million to 1.34 million (INETL, 2022), but it did not provide specific indicators for Tetun speakers.
3.2. Orthography
Tetun INL is based on the Latin alphabet, consisting of 5 vowels: a, e, i, o, u, and 21 consonants: b, d, f, g, h, j, k, l, ll, m, n, ñ, p, r, rr, s, t, ’, v, x, z (of Linguistics, INL). The letters c, q, w and y are not used in Tetun except for the proper names and international symbols. Accented vowels á, é, í, ó, ú are also used, and the apostrophe (’) denotes a glottal stop. Additionally, the hyphen is also used to indicate monosemantic compound words. Some basic phrases in Tetun are presented in Table 1.
Table 1. Examples of Basic Tetun Phrase.
Tetun English Tetun English
Dadeer di’ak! Good morning! Di’ak ka lae? How are you?
Ita-nia naran saida? What is your name? Ita-boot hela iha ne’ebé? Where do you live?
3.3. Morphology
Morphology is conventionally divided into inflection and word formation, with word formation further classified into derivation and compounding (Aronoff, 1983). Inflection refers to the different syntactic variations of a word that do not alter its core meaning, while word formation involves the creation of new nouns, verbs, and adjectives. Derivation creates a new word from an existing one, whereas compounding combines two or more words to form a new word.
Morphological processes such as circumfixes and reduplication also contribute to both the formation of new words and the modification of existing word structures. Circumfixes involve the simultaneous addition of a prefix and a suffix to a base word, while reduplication is a morphological process in which a part of a word is copied, either fully or partially, to form a new word that may have additional morphemes attached to it (Greksáková, 2018; Urbanczyk, 2017).
Tetun does not have rich inflectional and derivational morphology, with only a few inflectional affixes (Greksáková, 2018; Hull and Correia, 2005; van Klinken et al., 2002). Tetun affixes include both native Tetun elements and those derived from Portuguese. Prefixes are exclusive of native Tetun, whereas suffixes can derive from either native Tetun or Portuguese. In compounding, words are combined using hyphens, exclusively with native Tetun words. Additionally, Tetun uses circumfixes and reduplication within its native vocabulary and adopts zero derivation for Portuguese-derived words. Examples of Tetun inflection and derivation are provided in Table 2.
Table 2. Examples of Tetun Inflection and Derivation. Suffix dór is used in both native Tetun and Portuguese loanwords.
Prefixes Suffixes
Native Tetun Native Tetun Portuguese Loanwords
hadame (reconcile) susun (breast) selebrasaun (celebration)
nakfera (break) sala-na’in (sinner) ezatamente (exactly)
namkari (scatter) nakar-teen (naughty) doadores (donors)
hakbesik (get closer) hemudór (drinker) tokadór (musician)
In Tetun, both circumfixes and reduplication are not as widely used as in other languages. The circumfixes in Tetun are not productive (Greksáková, 2018) and are confined to simple verbs, typically consisting of one or two syllables derived from verbs (Hull and Correia, 2005). Reduplication is similarly limited, being applied only to nouns, adjectives, adverbs, and numerals, with only a few instances of its use for pluralization (Greksáková, 2018).
3.4. Portuguese Loanwords
A significant portion of Tetun’s verbs, nouns, and adjectives are derived from Portuguese, where this influence is particularly noticeable in the news media, such as newspapers (Hajek and van Klinken, 2019; Greksáková, 2018; van Klinken and Hajek, 2018; van Klinken et al., 2002). Klinken et al. (van Klinken et al., 2002) highlighted that the prevalence of Portuguese loanwords can be traced back to Portuguese-educated political leaders who continued to use Portuguese in their homes after 1975. As these leaders frequently appeared in the news media, the incorporation of Portuguese loanwords into Tetun rapidly increased.
Table 3. Example of Portuguese Loanwords.
Vebs Nouns Adjectives
estuda (estudar, study) serveja (cerveja, beer) baratu (barato, cheap)
kanta (cantar, sing) estudante (estudante, student) forte (forte, strong)
organiza (organizar, organize) eskola (escola, school) rápidu (rápido, fast)
Klinken and Hajek (van Klinken and Hajek, 2018) studied a selection of seven articles from different newspapers in 2009 and reported an average of 32% of words are Portuguese loanwords. Similarly, Greksáková (2018) highlighted 35% of Portuguese loanwords in the analysis of 73,892 words from interview transcripts. Moreover, Hajek and Klinken (Hajek and van Klinken, 2019) described Tetun’s influence from Portuguese in newspaper and technical writing, rising to over 40%, with headlines often almost entirely in Portuguese. In a recent study, de Jesus and Nunes (2024a) reported 28.20% of Portuguese loanwords in Tetun when analyzing approximately 10.69 million words extracted from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) for an interval time from 2017 to 2023. Additionally, they observed a 5.09 percentage point increase in Portuguese loanwords when comparing documents created before and after 2017 in the Labadain-30k+ dataset. Examples of Portuguese loanwords are presented in Table 3.
4. Dataset
In this work, we employed the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), comprising 33,550 Tetun documents acquired through web crawling. The dataset was thoroughly audited by native Tetun speakers at the document level and comprised a diverse range of categories, including news articles, Wikipedia entries, legal and government documents, and research papers, among others (de Jesus and Nunes, 2024a). A detailed description of the dataset is provided in Table 4, with a summary grouped by category of documents presented in Table 5. This dataset was employed to develop Labadain-Stopwords, Labadain-Stemmer, and Labadain-Avaliadór, which were subsequently used to evaluate the retrieval effectiveness of Tetun ad-hoc text retrieval.
Table 4. Description of the Labadain-30k+ Dataset. *Tokens consist of words and numbers.
Total documents in the dataset 33,550
Total paragraphs in the content 334,875
Total sentences in the content 414,370
Total tokens in the corpus 12,300,237
Vocabulary in the corpus 162,466
Table 5. Summary of the Labadain-30k+ Dataset.
Category #docs Proportion
News articles 30,150 89.87%
Wikipedia documents 1,455 4.34%
Legal/government documents 1,223 3.65%
Technical documents 211 0.63%
Blogs and Forums 145 0.43%
Advertisements/announcements 124 0.37%
Research papers 83 0.25%
Personal pages 74 0.22%
Institutional information 53 0.16%
Correspondence letters 32 0.10%
5. Methodology
To establish baselines for Tetun ad-hoc text retrieval, we employ the methodology illustrated in Figure 1. The process begins with the creation of a Tetun stopword list, continues with the development of a stemmer and a test collection, and concludes with experiments to establish the baselines. Each stage is described in the following subsections.
Figure 1. Methodology for Establishing Baselines in Tetun Ad-Hoc Text Retrieval.:
5.1. Labadain-Stopwords Construction
This initial stage focuses on constructing a Tetun stopword list. Since manually creating stopword lists is both time-consuming and expensive, we adopted a corpus-based approach using the Labadain-30k+ dataset. Candidate stopwords were generated using frequency- and network-based detection methods, and the resulting lists were merged and validated by two native Tetun speakers to produce the final Tetun stopword list (called Labadain-Stopwords).
Building on the findings of Ali et al. (2024), which demonstrated the superior effectiveness of network-based methods compared to traditional frequency-based techniques for stopword detection, we extended the evaluation to Tetun using the Labadain-Stopwords as the ground-truth set. We then compared the results with those for Portuguese and English to gain further insight.
5.2. Labadain-Stemmer Development
This stage focused on developing the stemmer algorithms for Tetun, called Labadain-Stemmer. Since a substantial portion of Tetun verbs, nouns, and adjectives are Portuguese loanwords, and Tetun suffixes encompass Portuguese-derived words and native Tetun, we created three stemmer variants: light, moderate, and heavy. The light variant removes only the suffixes of Portuguese loanwords, the moderate variant addresses both Portuguese loanwords and native Tetun suffixes, and the heavy variant handles Portuguese loanword suffixes, as well as native Tetun prefixes and suffixes.
To evaluate the proposed stemmer, we conducted both intrinsic and extrinsic assessments. For the intrinsic evaluation, we systematically extracted a subset of vocabularies from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) and collaborated with native Tetun-speaking students to construct a sample of words. These students assessed the sample word list provided to establish a ground truth list, with each word paired with its corresponding lemma (root). The ground truth set was then used to evaluate the accuracy of each stemmer variant using the Paice metrics (Paice, 1994). For extrinsic evaluation, we tested the effectiveness of Tetun stemmers in the ad-hoc text retrieval task.
5.3. Labadain-Avaliadór Building
Since no test collection exists for Tetun ad-hoc text retrieval, this stage focused on creating one following TREC guidelines. Native Tetun-speaking students developed queries by examining real-world search logs and the document collection sourced from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b). The same students also assessed the relevance of query-document pairs using a user-friendly interface we developed, with the document pooling process automated to streamline the assessment workflow. The resulting Tetun test collection is called Labadain-Avaliadór.
5.4. Tetun Ad-Hoc Retrieval Baseline Establishment
This stage focused on investigating various retrieval strategies for Tetun ad-hoc text retrieval. Documents and queries were initially preprocessed by converting text to lowercase, normalizing apostrophes, removing punctuation and special characters, tokenizing into individual tokens, and then performing document indexing, retrieval, and ranking to establish the baselines. Additional preprocessing steps, such as handling accented letters, apostrophes, and hyphens, were applied individually to assess their impact on retrieval effectiveness relative to the baselines. The process also included stopword removal and stemming.
The features that demonstrated improvements over the baseline were selected and combined for subsequent experiments to create the baselines. This approach was applied to both document titles (short text) and content (long text), employing various retrieval and ranking models. The effectiveness of these preprocessing steps and models was then assessed using various evaluation metrics to identify the most effective retrieval strategy for Tetun ad-hoc text retrieval.
6. Labadain-Stopwords Construction
This section introduces the frequency- and network-based approaches used to create Labadain-Stopwords, a Tetun stopword list. It describes the methodology for constructing the list and compares the effectiveness of network-based methods with traditional frequency-based techniques for stopword detection.
6.1. Overview
Frequency- and network-based approaches were applied in the development of the Labadain-Stopwords. Frequency-based methods, such as TF, IDF, and TF-IDF, rely on term-weighting techniques to identify frequently occurring words. TF measures the frequency of a term within a document, IDF evaluates the importance of a term by assessing how many documents in the collection it contains, and TF-IDF is the product of these metrics, representing the importance of a term within a document relative to its occurrence across the entire collection.
Network-based methods exploit the topological properties of co-occurrence networks modeled as directed graphs, including in-degree, out-degree, and degree. The in-degree represents the number of incoming connections, indicating how often a word is preceded by others. The out-degree captures the number of outgoing connections, showing how frequently a word precedes subsequent terms. The degree is defined as the sum of the in-degree and out-degree.
6.2. Approach
The Labadain-Stopwords was constructed using the Labadain-30k+ dataset (de Jesus and Nunes, 2024b). The process began with preprocessing steps, including lowercase, normalizing apostrophes, removing punctuation, special characters, numbers, and extra spaces, followed by tokenization using the Tetun tokenizer (de Jesus and Nunes, 2024e) and deduplication to create a vocabulary. The traditional term-weighting techniques (TF, IDF, and TF-IDF) were then applied to the vocabulary to assign weights to each word.
To analyze network properties, we constructed a vocabulary-level co-occurrence network as a directed graph from the preprocessed text, where each word corresponds to a node. For each node, we calculated in-degree (number of incoming links), out-degree (number of outgoing links), and degree (the sum of incoming and outgoing links), thereby quantifying the connectivity of words within the network.
Using these scores, the top 1,000 words from each method were selected in descending order based on their scores to create lists of potential stopwords. These lists were then merged, with duplicates and misspelled words excluded, to produce a candidate stopword list. Two native Tetun speakers—a Ph.D. student and an undergraduate student—reviewed and validated this list, resulting in Labadain-Stopwords, containing 160 Tetun stopwords (de Jesus and Nunes, 2025c). The complete list with English translations is provided in Appendix 12.1.
Some stopwords appeared in misspelled forms, such as for “ne’ebé” (meaning “which/that” in English) was found in variations like “nebe”“neebe”, and “neebé”. These variations were compiled into a separate list of stopword variations, which was subsequently used to develop a stopword corrector for application during the preprocessing step.
6.3. Experiment and Evaluation
In the network-based approach for detecting stopwords proposed by Ali et al. (2024), Tetun stopwords were manually translated from the English stopwords in NLTK [^7] https://www.nltk.org to establish the ground-truth set. In this study, we used the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) and evaluated the effectiveness of the approach with Labadain-Stopwords. To further assess performance across both low- and high-resource languages, we also conducted experiments with English and Portuguese.
For Portuguese and English, we used documents extracted from the CC-100 dataset (Wenzek et al., 2020) and employed stopword lists from NLTK as the ground truth. The process of assigning weights to Portuguese and English words followed the same approach used for Tetun. A summary of the datasets used to create the stopword lists is provided in Table 6.
Table 6. Summary of Tetun, Portuguese and English Datasets Used for Stopword Detection.
Description Tetun Portuguese English
Total number of documents 33,550 3,153 624
Total number of words 11,928,821 613,736 667,584
Total vocabulary size 146,783 45,860 31,390
For evaluation, we used precision at $n$ (P@$n$) to measure the proportion of stopwords among the top-$n$ words. While Ali et al. (2024) limited their analysis to P@200, we extended the P@$n$ cutoff to 1,000. For this purpose, we applied intervals of approximately 25 for cutoffs up to 100 and intervals of 250 for cutoffs between 100 and 1,000.
6.4. Results
The results of the experiment with Tetun are presented in Table 7, demonstrating that network-based approaches generally outperform traditional term weighting methods in identifying stopwords. Specifically, in-degree consistently demonstrates superior performance across most cutoffs, except at P@75, where degree slightly surpasses it. At P@10, P@25, and P@1000, in-degree and degree achieve identical performance scores. Notably, at the P@10 cutoff, all techniques perform equally well, achieving perfect precision. Among traditional term weighting methods, the results are comparable, with IDF slightly outperforming TF and TF-IDF at P@500 and P@1000.
Table 7. Stopword Precision for Tetun.
Approach P@10 P@25 P@50 P@75 P@100 P@250 P@500 P@750 P@1000
In-degree 1.0000 0.9600 0.8400 0.7200 0.7000 0.4720 0.3080 0.2347 0.1930
Out-degree 1.0000 0.8800 0.8000 0.6933 0.6000 0.4240 0.2900 0.2160 0.1780
Degree 1.0000 0.9600 0.8200 0.7333 0.6500 0.4640 0.3000 0.2253 0.1930
TF 1.0000 0.9200 0.6800 0.6000 0.5200 0.3600 0.2480 0.1933 0.1610
IDF 1.0000 0.9200 0.6800 0.6000 0.5200 0.3600 0.2540 0.1973 0.1640
TF-IDF 1.0000 0.9200 0.6800 0.5867 0.5100 0.3560 0.2500 0.1947 0.1620
When evaluated on Portuguese, similar patterns were observed, as shown in Table 8, with network-based methods again demonstrating superior performance. The degree slightly surpassed the in-degree at P@50, P@250, P@500, and P@750. At P@10, P@25, P@100, and P@1000, both in-degree and out-degree achieved identical scores. In-degree outperformed degree at P@75. At P@10 and P@25, all techniques performed equally well, achieving perfect precision. At P@1000, all methods achieved identical scores. Traditional term-weighting approaches yielded identical results across all evaluated cutoffs, though slight variations appeared at certain cutoffs when the dataset size was reduced.
Table 8. Stopword Precision for Portuguese.
Approach P@10 P@25 P@50 P@75 P@100 P@250 P@500 P@750 P@1000
In-degree 1.0000 1.0000 0.9000 0.7600 0.6700 0.3800 0.2260 0.1587 0.1260
Out-degree 1.0000 0.9600 0.9600 0.8000 0.6700 0.3560 0.2180 0.1573 0.1270
Degree 1.0000 1.0000 0.9400 0.8133 0.7100 0.3680 0.2240 0.1587 0.1240
TF 1.0000 0.9600 0.9000 0.7600 0.6500 0.3320 0.2000 0.1480 0.1200
IDF 1.0000 0.9600 0.9000 0.7600 0.6500 0.3320 0.2000 0.1480 0.1200
TF-IDF 1.0000 0.9600 0.9000 0.7600 0.6500 0.3320 0.2000 0.1480 0.1200
Similarly, in English, network-based approaches maintained their advantage, as shown in Table 9. The results mirrored those of Portuguese, with degree slightly outperforming in-degree at P@50, P@250, P@500, and P@750. In-degree and degree attained identical scores at P@100, while in-degree outperformed degree at P@75. As with the other languages, all methods achieved perfect precision at P@10 and P@25 and identical scores at P@1000. As in Portuguese, traditional term-weighting methods in English yielded identical results across all evaluated cutoffs.
Table 9. Stopword Precision for English.
Approach P@10 P@25 P@50 P@75 P@100 P@250 P@500 P@750 P@1000
In-degree 1.0000 1.0000 0.9400 0.9333 0.8200 0.4120 0.2280 0.1573 0.1220
Out-degree 1.0000 1.0000 0.9800 0.8400 0.7600 0.4160 0.2280 0.1573 0.1220
Degree 1.0000 1.0000 0.9800 0.9067 0.8200 0.4280 0.2300 0.1613 0.1220
TF 1.0000 1.0000 0.9600 0.8933 0.7500 0.4240 0.2280 0.1573 0.1220
IDF 1.0000 1.0000 0.9600 0.8933 0.7500 0.4240 0.2280 0.1573 0.1220
TF-IDF 1.0000 1.0000 0.9600 0.8933 0.7500 0.4240 0.2280 0.1573 0.1220
6.5. Discussion
To examine the stopword detection approaches across different levels, we divide precision into lower cutoffs (up to P@25), mid-range cutoffs (P@50 to P@100), and higher cutoffs (P@250 to P@750). At lower cutoffs, all methods yielded similar results, with network-based approaches, such as in-degree and degree, slightly outperforming traditional term-weighting methods by a small margin of +0.04 points in Tetun and Portuguese at P@25.
In mid-range cutoffs, network-based methods maintained their advantage, surpassing traditional methods by up to +0.06 points in Portuguese, +0.07 points in English, and +0.18 points in Tetun. Among network-based methods, degree consistently outperformed in-degree at these mid-range cutoffs.
In higher cutoffs, network-based methods still outperformed traditional term weighting approaches, with in-degree consistently delivering the best results for Tetun and Portuguese, showing improvements of up to +0.18 points. However, in English, the degree marginally surpassed the in-degree. These results indicate that network-based methods maintain a stronger advantage as the stopword list expands, except in English at P@1000, where all methods produced identical scores, likely due to characteristics of the dataset.
Overall, in Tetun, in-degree was slightly more effective than degree, while in English, degree marginally outperformed in-degree. In Portuguese, both methods performed similarly (see Figure 2). Since the degree is defined as the sum of in-degree and out-degree and yields performance comparable to in-degree, the latter offers advantages in computational efficiency.
Figure 2. Comparison of the Network-Based Approach Performance at Mid-Range and Higher Cutoff Levels.:
For traditional term-weighting approaches, IDF outperformed TF and TF-IDF at higher cutoffs in Tetun. At mid-range cutoffs, IDF and TF achieved identical scores, while at lower cutoffs, TF, IDF, and TF-IDF all produced identical scores. In contrast, for Portuguese and English, TF, IDF, and TF-IDF yielded identical results across all cutoffs. This consistency in Portuguese and English may be due to the more structured and mature linguistic resources available for these languages, such as well-established stopword lists and corpora, which minimize variations in term weighting effectiveness. In Tetun, the language’s lesser-resourced nature likely results in greater sensitivity to different weighting methods, leading to performance differences at higher cutoffs. Furthermore, the total number of stopwords we developed for Tetun is comparable to other LRLs such as Marathi (99 stopwords), Bengali (114 stopwords), and Hindi (165 stopwords) (Sahu et al., 2023); and Kinyarwanda (80 stopwords) and Kirundi (59 stopwords) (Niyongabo et al., 2020).
6.6. Conclusion
This study highlights the superiority of network-based approaches, particularly in-degree and degree, over traditional term weighting methods for stopword detection in both high- and low-resource languages, especially when dealing with larger stopword sets. Although traditional term weighting and network-based methods perform comparably at smaller cutoffs (up to 25 terms), network-based approaches demonstrate greater effectiveness as the number of evaluated terms increases. For smaller stopword lists, the differences between methods are less significant. However, when working with lists of 25 or more stopwords, network-based approaches are recommended for their superior performance at mid-range and higher cutoffs. The consistent in-degree performance observed in Tetun is aligned with the findings reported by Ali et al. (2024), further validating the effectiveness of network-based methods for stopword detection tasks, specifically in under-resourced scenarios.
7. Labadain-Stemmer Development
This section describes the development of Labadain-Stemmer, a stemming algorithm specifically designed for Tetun. It covers the identification of Tetun affixes and the creation of Labadain-Stemmer variants tailored to the language. Additionally, it details the process of generating a sample of words, which native Tetun speakers assessed to serve as the ground truth for evaluating the accuracy of Labadain-Stemmer. Finally, the section presents the experiments conducted, their results, and corresponding discussion, concluding with a summary of limitations and key observations.
7.1. Tetun Affixes
This study focuses on commonly used affixes in Tetun (van Klinken et al., 2002; Greksáková, 2018; Hull and Correia, 2005), excluding circumfixes and reduplication due to their limited productivity and usage, as discussed in Subsection 3.3. The native Tetun prefixes are “ha”“nak”, and “nam”, while the native suffixes comprises “n”“-nain”“-teen”, and “dór”. Additionally, Portuguese-derived suffixes are adapted from the list of Portuguese suffixes used in the Portuguese stemmer in Snowball (Snowball, 2005a), as presented in Table 30 in the Appendix 12.2. Since Tetun has few inflectional affixes, stemming native Tetun words is a straightforward process that involves matching words with a predefined list of Tetun prefixes and suffixes at the beginning and end of each word.
7.2. Stemmer Variants
Tetun consists of both native words and a significant number of Portuguese loanwords, particularly verbs, nouns, and adjectives (Hajek and van Klinken, 2019; Greksáková, 2018; van Klinken and Hajek, 2018; van Klinken et al., 2002). To address this linguistic mix, the Labadain-Stemmer is designed with three variants: lightmoderate, and heavy. Each variant is detailed in the following subsections.
7.2.1. Light Stemmer
The light stemmer is designed to remove suffixes from Portuguese-derived words used in Tetun. This variant adapts the Portuguese stemmer from Snowball, incorporating a customized list of Portuguese suffixes. These suffixes were modified based on the loanword transformation rules defined by the INL (of Linguistics, INL), as detailed in Table 10.
Table 10. Rules for Transforming Portuguese-Derived Words into Tetun Which Applied in Suffix Transformations.
Rule (Portuguese $\rightarrow$ Tetun) Effect on Suffix Example (Portuguese $\rightarrow$ Tetun)
ão $\rightarrow$ aun asaun comemoração $\rightarrow$ komemorasaun (celebration)
ss, c (before e, i), ç (before a, o, u) $\rightarrow$ s saun discussão $\rightarrow$ diskusaun (discussion)
qu, c (before a, o, u) $\rightarrow$ k ik + amente automaticamente $\rightarrow$ automatikamente (automatically)
g (before e, i) $\rightarrow$ j lojia tecnologia $\rightarrow$ teknolojia (technology)
s (between vowels) $\rightarrow$ z oza poderosa $\rightarrow$ poderoza (powerful)
ê $\rightarrow$ é énsia competência $\rightarrow$ kompeténsia (competence)
â $\rightarrow$ á ánsia ignorância $\rightarrow$ ignoránsia (ignorance)
o $\rightarrow$ u u infermeiro $\rightarrow$ infermeiru (nurse)
The Tetun light stemmer is a simplified adaptation of the Portuguese stemmer, designed to handle loanwords while accounting for Tetun’s unique morphological characteristics. It retains the linguistic regions utilized in the original Portuguese stemmer, which were adapted from the Spanish stemmer in Snowball (Snowball, 2005b). The definitions of these linguistic regions, as applied in the Tetun light stemmer algorithm, are provided in Table 29 of Subsection 12.2.
The Tetun light stemmer processes words sequentially using a list of suffixes developed to account for the specific features of Portuguese loanwords in Tetun (see Table 30). The stemming procedure is summarized below, with the corresponding algorithm provided in Algorithm 12.2:
  1. (1)
    Word length validation: After receiving an input word, the algorithm begins by validating its length. If the word contains fewer than four characters, it is returned without stemming.
  2. (2)
    Standard suffix removal: For words longer than three characters, the algorithm searches for the longest matching suffix from the general suffix list. If a matching suffix is found in a specific region of the word, the suffix is deleted or replaced accordingly.
  3. (3)
    Verb suffix removal: If no suffix is removed in step (2), the algorithm checks for verb-specific suffixes. It is removed if a matching suffix is found within the appropriate region of the word.
  4. (4)
    Residual suffix removal: If neither of the previous steps results in suffix removal, the algorithm looks at the remaining simple suffixes list and removes it as the final step.
  5. (5)
    Return original word: If none of the steps result in suffix removal, the input word is returned unchanged.
7.2.2. Moderate Stemmer
The moderate stemmer extends the functionality of the light stemmer by handling suffixes from both Portuguese loanwords and native Tetun. It adheres to the same algorithm as the light stemmer (outlined in Algorithm 12.2), with the addition of a new step of 4.1, specifically designed to process native Tetun suffixes. This additional step is executed between steps 4 and 5 of the algorithm.
7.2.3. Heavy Stemmer
The heavy stemmer builds on the functionality of the moderate stemmer by introducing the removal of native Tetun prefixes. In this variant, the processing of native Tetun prefixes is integrated between steps 4.1 and 5 of the algorithm.
7.3. Text Sample Construction for Evaluation
This subsection outlines the creation of sample words that are used for intrinsic experimental and evaluation purposes. Selecting sample words to assess stemming performance poses challenges due to potential bias and limited generalization. To mitigate this, we designed a systematic methodology to construct this sample from a dataset containing a diverse collection of categories and sources, with human involvement in the loop.
Figure 3. Process of Constructing a Text Sample for Evaluating Tetun Stemmer’s Performance.:
The process of constructing sample words used for the experiment and evaluation is illustrated in Figure 3. First, we preprocessed the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), which involved lowercase, normalizing apostrophes, and removing punctuation, special characters, numbers, and extra spaces. After this, the text was tokenized into individual words using the Tetun tokenizer (de Jesus and Nunes, 2024e), and deduplication was performed to remove duplicate words. This preprocessed text was then tokenized into individual words using the Tetun Word Tokenizer (de Jesus and Nunes, 2024d), and deduplication was performed to remove duplicate tokens. Stopwords were normalized and subsequently excluded, along with tokens shorter than four characters. To further refine the candidate sample words, the Tetun LID model (de Jesus and Nunes, 2024c) was applied with a threshold score of 0.95 to filter out words that did not meet this criterion. Finally, the refined candidate sample words were validated by native Tetun speakers to produce the final set of sample words.
The sample word verification with human involvement was carried out by six native Tetun speakers, consisting of one Ph.D. and five undergraduate students. Each student was tasked with verifying approximately 2,732 words, sorted in ascending alphabetical order. They checked the correctness of each word using both the INL dictionary (Correia et al., 2005) and the Portuguese loanword dictionary (Greksáková, 2018) as reference materials. During the verification process, a considerable number of misspelled words were identified, such as the word “konsiderasaun” (consideration, in English) appearing as [ konsidersaunkonsiderasaunnkonsideransaun ]. Additionally, some words originating from Tetun Terik, Tetun DIT, or other variants not present in the reference dictionaries were excluded from the final sample. A summary of the resulting sample of words from this process is presented in Table 11.
Table 11. Summary of Words Produced at Each Stage of the Sample Word Generation Process. Approximately 88.4% reduction in the number of words was observed after applying the LID model, attributed to the model’s average score per word of 0.3943, with a threshold set to 0.95.
Description Total of Words
Initial total number of words 146,387
Remaining words after removing stopwords 146,204
Remaining words after excluding words with fewer than four characters 141,487
Remaining words after applying LID model 16,391
Remaining words after manual verification by human assessors 1,839
7.4. Ground Truth Development
The development of the ground truth set for evaluating the Labadain-Stemmer performance was carried out by the same six Timorese students. To familiarize the assessors with the evaluation process, five example pairs from the sample of 1,839 words (see Table 11) were provided during the training session. These pairs included the original words and their corresponding stemmed forms generated by each Labadain-Stemmer variant. After this initial training, the complete list of input words and their stemmed results was distributed to two students per stemmer variant for evaluation. Their primary task was to determine whether each word correctly stemmed to its root form. When a word was incorrectly stemmed, the students provided the correct root form, using the suffixes detailed in Table 30 for Portuguese-derived words and the Tetun affixes described in Subsection 7.1 to guide their decisions.
Inter-annotator agreement was calculated to ensure consistency and reliability among the annotators. Discrepancies between annotators were analyzed and discussed, allowing them to reach a consensus for each stemmer variant. This procedure was followed by all annotators during the evaluation of the different stemmer variants. Inter-annotator agreement was measured using Cohen’s kappa, as presented in Table 12. In the final stage, the annotators pooled their evaluations and resolved any remaining discrepancies to finalize the correct stemmed forms. These consensus-based results were then used to compile the ground truth set, which is summarized in Table 13.
Table 12. Cohen’s Kappa Score for Inter-Annotator Agreement in the Construction of the Ground Truth Set.
Algorithm k-Score
Light stemmer 0.7006
Moderate stemmer 0.6990
Heavy stemmer 0.7683
Table 13. Summary of the Ground Truth Set and Word Length.
Description All Words Portuguese Loanwords Native Tetun Words
Total number of words 1,839 81.79% 18.21%
Minimum character count per word 4.00 4.00 4.00
Maximum character count per word 20.00 20.00 15.00
Average character count per word 9.50 10.21 5.97
7.5. Intrinsic Evaluation
In intrinsic evaluation, the Paice metric (Paice, 1994) was used to evaluate the stemmer variant quality by measuring how effectively they reduce various word forms to a common root. This metric balances understemming and overstemming effects, both of which impact precision and recall in text-processing tasks. In IR, a high understemming lowers recall, resulting in relevant documents not being retrieved, while a high overstemming hurts precision by retrieving many irrelevant documents.
Paice introduced four intrinsic methods to assess stemming performance: understemming index (UI), overstemming index (OI), stemming weight (SW), and error rate relative to truncation (ERRT). UI measures how often the stemmer fails to reduce related words to the same root, while OI calculates the frequency of incorrectly merging unrelated words into the same root. SW is the ratio of OI/UI, representing the trade-off between overstemming and understemming. A lower value of SW indicates more understemming, whereas a higher value suggests a tendency toward overstemming. ERRT evaluates the stemmer’s ability to balance understemming and overstemming. This involves computing UI and OI values for various truncation lengths to establish a truncation line, which serves as a baseline for stemmer performance. Any reasonable stemmer should have its (UI, OI) point located between the truncation line and the origin, with better performance indicated by a position further away from the truncation line or closer to the origin.
7.6. Experimental Setting
To compute the Paice metric, a list of words is first organized into conceptual groups based on semantic and morphological relationships. These groups serve as the target, and an ideal stemmer should conflate words according to these conceptual groupings. The stemmers were then applied to the word list, and their performance was evaluated by measuring how accurately they matched the predefined conceptual groups. Examples of these conceptual groupings are provided below, where the root word is shown on the left side and its corresponding conflated words are listed on the right side, separated by a colon delimiter.
’ajente’: [’ajénsia’, ’ajénsias’]
    ’akompañ’: [’akompaña’, ’akompañadu’, ’akompañamentu’, ’akompañante’]
    ’akontes’: [’akontese’, ’akontesimentu’, ’akontesimentus’]
    ’hatete’: [’hatete’, ‘hateten’]
    ’kbiit’: [’kbiit-laek’, ’kbiit-na’in’, ’kbiit’]
    ’komunik’: [’komunikadu’, ’komunikadus’, ’komunikadór’, ’komunikasaun’, ’komunikativa’]
    ’otél’: [’otél’]
The application of the stemmer to conceptual groups resulted in understemming, overstemming, and the relative accuracy of the stemmers, represented by ERRT. To calculate ERRT, a baseline was established using length truncation, where the words in the list were truncated to their first $n$ letters, with $n$ set to 7, 8, and 9. The overstemming and understemming measures of these truncated lists define the truncation line.
7.7. Evaluation and Results
Using the Paice metric, we calculated the ERRT value for each stemmer variant by drawing a line from the origin through the point representing its understemming and overstemming indexes (UI, OI) and extending it to intersect the truncation line. The ERRT is calculated by dividing the distance from the origin to the (UI, OI) point by the distance from the origin to the truncation line intersection. An ideal stemmer variant has low UI and OI values, indicating better performance when positioned closer to the origin or further away from the truncation line. Figure 4 presents the UI and OI values for each stemmer with the truncation line, showing that the heavy and moderate stemmer variants slightly outperformed the light variant.
Figure 4. UI vs. OI Plot Showing ERRT Distances.:
The UI, OI, SW, and ERRT values are presented in Table 14. As expected, the light variant exhibits the highest understemming value (by +0.007 points), while the moderate and heavy variants have identical lowest ERRT values (both lower by approximately -0.009 points), and all variants have the same overstemming values. Further investigation revealed that the difference in words stemmed from the heavy variant compared to the moderate variant was limited to only six words. This is due to the small proportion of native Tetun words, which make up only 18.21% of the total (see Table 13), and the limited number of Tetun prefixes (outlined in Subsection 7.1).
Table 14. Analysis of the Stemming Algorithms’ Performance Using Paice Metrics.
UI OI SW ERRT
Light 0.312062 0.000017 0.000056 0.481367
Moderate 0.305049 0.000017 0.000057 0.472802
Heavy 0.305049 0.000017 0.000057 0.472802
Regarding the overstemming value, since native Tetun has few inflectional forms in the word list, applying both the moderate and heavy variants had no impact on the overstemming value. The minimal difference in the ERRT values presented in Table 14 indicates that the lightmoderate, and heavy stemmer variants perform quite similarly.
7.8. Discussion
Tetun, as a language with relatively few inflectional affixes (Greksáková, 2018; Hull and Correia, 2005; van Klinken et al., 2002), often includes short affixes, such as the prefix “ha” and the suffix “n”, which present challenges to stemming algorithms in correctly handling native Tetun affixes. Some verbs and nouns begin or end with these characters, though they are not affixes, as seen in words like “halimar” (play), “hariis” (bathe), “aman (father), “inan (mother), “ibun (mouth), “liman (hand), “ulun (head), among others. Additionally, removing these characters from certain words changes their meaning. For example, removing “ha” from “halimar” (play) results in “limar” (rasp), and removing the suffix “n” from “liman (hand) becomes “lima” (five).
Given that three variants of the Labadain-Stemmer show similar performance, it might be affected by the characteristics of the sample of words used for evaluation. Factors such as the proportion of native Tetun words, the presence of affixes, and the term distribution within the sample could influence the stemming algorithms’ outcomes. Developing a more balanced sample of word composition could provide deeper insight into the observed results.
7.9. Limitations
The Instituto Nacional de Linguística (INL) launched the Kursu Gramátika Tetun (Tetun Grammar Course) in 2005, which served as a reference for teachers, translators, journalists, and students (Hull and Correia, 2005). It includes several Tetun affixes, such as prefixes [ hak, na, ma ] and suffixes [ -laek, k ]. However, more recent research by  Greksáková (2018) in 2018 reported that many of these affixes have been largely replaced by words such as “sai” (meaning “become”) and “laiha” (meaning “without”), making these affixes less productive in Tetun. As the INL has not updated its 2005 publication on Tetun grammar, the current state of Tetun morphology remains unclear.
Furthermore, the absence of linguistic experts in this study, due to the lack of funding to hire linguists, represents a limitation. Nevertheless, we have established a baseline that can serve as a foundation for future research in Tetun.
7.10. Conclusion
This study developed and assessed the effectiveness of the Labadain-Stemmer, incorporating suffixes of Portuguese loanwords and the affixes of native Tetun words. The Tetun affixes used were based on those commonly reported by Klinken et al. (van Klinken et al., 2002), the INL (Hull and Correia, 2005), and Greksáková (2018). To evaluate stemmer performance, we systematically constructed sample words and established a baseline for Labadain-Stemmer, testing three variants (light, moderate, and heavy). Results showed that integrating native Tetun affixes into the stemming process was marginally more effective than focusing solely on the suffixes of Portuguese loanwords.
However, one of the limitations in this study is the unbalanced representation of Portuguese loanwords and native Tetun words in the sample set. Future research should address this by using more balanced datasets that adequately represent both Portuguese loanwords and native Tetun words. Furthermore, the involvement of expert linguists specializing in Tetun will be crucial for enhancing the accuracy and overall effectiveness of the Labadain-Stemmer. To enable reproducibility, the stemmer algorithms have been released under the MIT License (de Jesus and Nunes, 2025b), to encourage further research and development in Tetun information retrieval.
8. Labadain-Avaliadór Building
This section provides an overview of the test collection and details the process of constructing Labadain-Avaliadór (avaliadór, a Tetun word meaning “evaluator”), a Tetun test collection for evaluation. It covers the dataset used, query formulation, document pooling, and relevance judgments.
8.1. Overview
The effectiveness of information retrieval systems relies on the availability of reliable test collections for evaluation. The traditional approach to building such collections follows the Cranfield paradigm (Cleverdon, 1967), which is widely adopted through the TREC evaluation campaigns (Harman, 1992). A TREC-style test collection typically comprises three core components: a document collection, a set of information needs (or topics), and relevance judgments.
8.2. Documents
The document collection comprises 33,550 Tetun documents sourced from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), each enriched with metadata such as title, URL, source, publication date, and content. This dataset was collected from web crawling and covers a wide range of categories, including news articles, Wikipedia entries, legal and government documents, research papers, technical documents, blogs, forums, and more (de Jesus and Nunes, 2024a). The diversity of its sources and topics makes this dataset particularly suitable for constructing a test collection for Tetun. A sample of the documents, formatted according to TREC guidelines, is shown in Figure 5. The collection is 84 MB in size, with approximately 12.3 million tokens and 162,466 unique tokens. A summary of the collection is provided in Table 15, and the length distribution of titles and content is illustrated in Figure 6.
Figure 5. Sample of Document Formatted Following TREC Guidelines: Original (left) and English translation (right).:
Table 15. Summary of Document Collection. Tokens comprise words and numbers, excluding punctuation and special characters.
Description Total Min Max Avg Std
Number of tokens (titles) 306,840 1 29 9.15 3.05
Number of tokens (content) 11,997,420 2 27,166 357.48 473.99
Figure 6. Length Distribution of Titles and Body Content in the Labadain-30k+ Dataset. In the right-hand figure, the x-axis of the document length distribution is limited to 1,000 tokens (words and numbers) for improved visualization.:
8.3. Query Formulation
Queries were collected from two sources: Google Search Console [^8] https://search.google.com/search-console for Timor News and the user search logs from the Timor News platform. [^9] https://www.timornews.tl The Google Search Console queries cover the period from November 1, 2021, to January 31, 2024, while the search logs from Timor News span from May 7, 2021, to January 31, 2024. Timor News is an online news agency based in Dili, Timor-Leste, founded in May 2019 and launched its news portal on May 7, 2019. The platform registered an average of 1,400 unique visitors per day and exclusively publishes news in Tetun.
The collected queries were compiled and distributed among five second-year undergraduate volunteers, all native Tetun speakers from Timor-Leste. The group comprised two students from Environmental Engineering, two from Information Systems, and one from Medicine. These students were tasked with developing queries following the established guidelines. Initially, each student was assigned 250 queries, resulting in a total of 1,250 queries analyzed. To understand user information needs, the students reviewed the provided search logs and either retained, modified, or formulated new queries based on the contextual information interpreted from the logs (see examples in Table 16).
Table 16. Examples of original query logs and their reformulations. Words highlighted in green background indicate newly added terms, while word in orange background indicates orthography correction.
Original query log Reformulated query
Problema lixu (Waste problem) Problema lixu iha Dili (Waste problem in Dili)
Soe bebe (Baby abandonment) Kazu soe bebé (Case of baby abandonment)
Konsumu tabaku (Tobacco consumption) Dadus konsumu tabaku (Tobacco consumption data)
Before beginning query development, the students attended a training session that provided practical examples of query formulation. Following this, three pilot testing sessions were conducted, during which each student created a query, defined the associated information need, and described the types of documents they would consider relevant.
The queries were then entered into a search prototype [^10] https://www.labadain.tl built on top of Apache Solr [^11] https://solr.apache.org using the BM25 ranking model. This prototype allowed the students to analyze the documents retrieved for each input query. For each query, each student selected four documents from the top 50 retrieved list, ensuring that one document represented each category: non-relevant, marginally relevant, relevant, or highly relevant. These sessions facilitated discussions about the results, highlighted challenges faced, and provided feedback to deepen their understanding of query development.
Figure 7. Sample of Topic Formatted According to TREC Guidelines: Original (left) and English translation (right).:
Subsequently, students were tasked to develop short queries following TREC best practices (Sanderson, 2010), ranging from three to five words, specifying information needs and describing the types of documents they expected the system to retrieve. Using the search prototype mentioned earlier, they input their queries and analyzed the retrieved documents. To finalize each query, they ensured that at least five relevant documents were identified. In total, 61 queries were developed. A sample query (or topic), formatted according to TREC guidelines, is shown in Figure 7. A summary of the queries is provided in Table 17, and their distribution across categories is shown in Figure 8. For categorization, we adapted the query topic categorization frameworks of Beitzel et al. (2004) and Rohatgi et al. (2021).
Table 17. Summary of Queries.
Description Value
Total of queries 61
Total number of three-word queries 37
Total number of four-word queries 22
Total number of five-word queries 2
Average numbers of words per query 3.43
Figure 8. Distribution of Queries Over Categories.:
8.4. Document Pooling
Since Query-document relevance judgments are carried out by human assessors, it is not feasible to evaluate every document in a large collection. To address this challenge, Spärck-Jones and van Rijsbergen (1975) introduced the pooling technique, in which a small subset containing a sufficiently representative sample of relevant documents is selected from the larger collection and provided to human assessors for relevance judgments (Sanderson, 2010).
Given the limited availability of retrieval models and techniques for LRLs, and to maximize the retrieval of relevant documents for constructing a robust test collection, we created the document pool using two retrieval models: BM25 and a language model (LM) with Dirichlet smoothing. BM25 is widely recognized for its effectiveness in ad-hoc retrieval (Robertson and Zaragoza, 2009), while the LM with Dirichlet smoothing has been shown to perform particularly well on short queries (Zhai and Lafferty, 2001). To balance the contributions of these models, we applied the balanced interleaving technique (Radlinski et al., 2008) to merge their results into a pool, which was then presented to assessors for evaluation.
Documents were indexed in separate instances of Solr, each configured with either the BM25 or the Dirichlet LM ranking models. The document retrieval and pooling process was fully automated and integrated with a relevance assessment interface to streamline the workflow. When a query was received, the system retrieved candidate documents ranked by each of the two models, merged the results into a pool, and presented the top 100 documents to the assessors for relevance judgments. The architecture of the retrieval system used for these assessments is shown in Figure 9.
Figure 9. General Architecture of the Retrieval System used for Relevance Judgments. The document highlighted in red indicates a duplicate that was excluded from the final list.:
8.5. Relevance Judgment
Five native Tetun-speaking students who developed the queries conducted relevance judgments for the query-document pairs, categorizing them into four graded levels of topical relevance: non-relevant, marginally relevant, relevant, and highly relevant, as proposed by Sormunen (2002), following the guidelines outlined in Subsection 8.3. A user-friendly web interface was created to streamline the process, allowing assessors to log in with individual accounts to conduct assessments. The interface used for the evaluation is illustrated in Figure 10.
Figure 10. Web Interface Used by Human Assessors for Conducting Relevance Assessments.:
Once logged in, the assessors began by selecting a query (label 1 in Figure 10) and reviewing the associated information needs and relevance criteria (label 2). They then evaluated each of the 100 documents, assigning a relevance score to each, and submitted their judgments (label 3). Upon submission, the system redirected the assessors to the homepage, displaying a list of queries. The option to reassess previously evaluated queries was automatically disabled.
Each assessor evaluated the same set of 61 queries, with 100 documents to assess per query, resulting in a total of 6,100 documents being assessed. Assessors were instructed to focus on topical relevance, check document details if the query was too long and not fully displayed on the interface, and disregard the retrieval order when determining the relevance of the document for the given query. The assessment process was completed within eight hours, and the number of queries assessed by each annotator per hour is illustrated in Figure 11. Documents not included in the judgment list were considered non-relevant.
Figure 11. Hourly Statistics of the Total Number of Queries Evaluated by Each Annotator.:
To assess inter-annotator reliability, we used Cohen’s kappa measure (Cohen, 1960), interpreting the strength of agreement according to the scale provided by Landis and Koch (1977). The overall average score of inter-annotator agreement among the five annotators is 0.4236, indicating moderate agreement, with detailed results presented in Table 18.
Table 18. Cohen’s Kappa scores for Inter-Annotator Agreement Among the Five Assessors.
HA1 HA2 HA3 HA4 HA5
Human annotator 1 (HA1) 0.4344 0.4999 0.4434 0.4380
Human annotator 2 (HA2) 0.4344 0.3745 0.3310 0.3500
Human annotator 3 (HA3) 0.4999 0.3745 0.4646 0.4199
Human annotator 4 (HA4) 0.4434 0.3310 0.4646 0.4063
Human annotator 5 (HA5) 0.4380 0.3500 0.4199 0.4063
Average kappa score 0.4236
Since all assessors evaluated the same queries, a majority voting approach was applied to determine each document’s relevance to its corresponding query, based on the scores assigned to the query-document pairs. According to the majority voting rule, the most frequently chosen label for each query-document pair must exceed 50% of the total votes to qualify as the majority label (Snow et al., 2008), meaning that at least three annotators must select the same label. The evaluation results revealed that, out of 6,100 documents assessed, 9.87% did not meet this threshold or resulted in ties (e.g., two groups of annotators selected different labels, such as annotators 1 and 2 choosing score 3, annotators 3 and 4 choosing score 1, and the fifth annotator selecting score 0). Details of these ties are presented in Table 19.
Table 19. Summary of Documents with Tied Scores.
Description Value
Total number of tied documents 602
Proportion of tied documents 9.87%
Minimum number of tied documents per query 1
Maximum number of tied documents per query 38
Average number of tied documents per query 10.20
Standard deviation in the number of tied documents per query 7.17
An approach to addressing tied scores is to use the tie-breaker strategy, which suggests that using a strong signal to break ties is more effective than a weak one (Wu and Fang, 2013). By applying this tie-breaker method to all instances of tied scores, we obtained the results shown in Table 20 (referring to the 1st round). However, after conducting an in-depth analysis of the tied scores, we observed significant discrepancies in some cases, such as ties between scores of 0 and 2 or 1 and 3. To resolve these inconsistencies, we re-invited three of the five original assessors for a second round of evaluations on the tied documents. During this phase, assessors were presented with the two tied score options from the initial assessment. The reassessment was conducted using Microsoft Excel, with separate tabs for each query and its corresponding documents, and was completed in approximately one hour and 15 minutes.
Table 20. Details of the Human Judgment Results.
Relevance Level 1st Round 2nd Round
3 - Highly relevant 710 11.64 566 9.59
2 - Relevant 1,102 18.07 1,054 17.86
1 - Marginally relevant 476 7.80 549 9.31
0 - Irrelevant 3,812 62.49 3,731 63.24
After completing the second round, we merged the evaluation results with those from the first round and applied a majority voting method, selecting the most frequent score for each query-document pair as the final relevance score. To ensure reliability, queries with 100 or more relevant documents or fewer than ten relevant documents were excluded (Sanderson, 2010; AleAhmad et al., 2009), resulting in the exclusion of two queries with ten or fewer relevant documents in the second round.
Table 21. Summary of the Final Test Collection. *Relevant documents consist of marginally relevant, relevant, and highly relevant.
Description Value
Total number of topics 59
Total number of qrels 5,900
Minimum number of relevant documents per query 11
Maximum number of relevant documents per query 99
Average number of relevant documents per query 36.76
Standard deviation of relevant documents per query 20.89
The final test collection, called Labadain-Avaliadór (de Jesus and Nunes, 2025a), contains an average of 36.76 relevant documents per query, detailed in Table 21—comprising 9.59% highly relevant, 17.86% relevant, 9.31% marginally relevant, and 63.24% non-relevant documents—as shown in the 2nd round column in Table 20. The distribution of document relevance per query is illustrated in Figure 12.
Figure 12. Total Number of Relevant and Non-relevant Documents per Query. Relevant documents consist of marginally relevant, relevant, and highly relevant.:
8.6. Results and Discussion
Table 20 shows that, after reassessing the tied documents from the first round results using the tie-breaker strategy, the number of highly relevant documents decreased by 144 and relevant documents by 48, while marginally relevant documents increased by 73, and non-relevant documents decreased by 81. This shift suggests that some documents initially classified as highly relevant or relevant were reclassified as marginally relevant, and some marginally relevant documents were reclassified as non-relevant. Additionally, 200 documents from two excluded queries were removed, indicating a change in relevance interpretation after resolving the ties. This reclassification adjusted the distribution of documents across categories, enhancing the overall quality of the test collection.
When analyzing the changes in document relevance, annotator 2 exhibited a distinct pattern, as illustrated in Figure 11. This annotator judged seven queries (700 documents) in the first hour at an average rate of 5.14 seconds per query-document pair, significantly faster than the average of approximately 8.18 seconds per pair observed among other annotators, who assessed an average of 4.4 queries (440 documents) in the first hour. This fast pace persisted until the seventh hour, leaving only two queries for the eighth hour. Additionally, Table 18 shows that annotator 2 had only fair agreement with three other annotators (2, 3, and 4), which contributed to increased discrepancies. These observations suggest that the quality of annotations from annotator 2 was lower, probably due to the speed of the evaluation process, which affected the changes in the relevance of the document observed after the second round of evaluations.
Table 22. Comparison of the Labadain-Avaliadór With Other LRLs.
Language #Docs #Topics Avg. Relevant Docs
Tetun 33,550 59 36.76
Sanskrit (Sahu and Pal, 2023) 7,057 50 8.54
Chichewa (Chavula and Suleman, 2021) 9,380 129 19
Citumbuka (Chavula and Suleman, 2021) 2,258 129 17
Cinyanja (Chavula and Suleman, 2021) 173 129 15
Hamshahri (AleAhmad et al., 2009) 166,774 65 36.18
Finally, Table 22 compares the Labadain-Avaliadór to other LRL collections. The Labadain-Avaliadór offers a balanced combination of scale and relevance density, making it a valuable resource for ad-hoc text retrieval in low-resource settings like Tetun. With 59 topics and 33,550 documents, it provides a moderately sized corpus that surpasses smaller collections like Sanskrit and Citumbuka but is smaller than larger collections like Hamshahri. Additionally, the diversity of its topics (see Figure 8) ensures its suitability for ad-hoc retrieval tasks.
The Labadain-Avaliadór has high relevance density with an average of 36.76 relevant documents per topic, the highest among the collections analyzed. This level of relevance provides a robust foundation for retrieval experiments, supported by a substantial pool of annotated relevance judgments that enable precise and reliable evaluations. Compared to collections with size variations of up to 50,000 documents, such as Sanskrit and Chichewa, the Labadain-Avaliadór stands out as a well-annotated collection, particularly suited for ad-hoc text retrieval in LRL contexts.
8.7. Conclusion
This study describes the development of Labadain-Avaliadór following TREC guidelines. Five native Tetun-speaking students conducted both the query development and the query-document relevance assessment. The queries were derived from real-world search logs in Tetun, and the test collection was graded on a scale from zero to four. The results indicate that the assessors agreed on the relevance of more than 90% of the query-document pairs assessments, with an average inter-annotator agreement of Cohen’s kappa score of 0.4236, indicating moderate agreement.
Approximately 10% of the 6,100 query-document pairs showed disagreement, resulting in tied scores. To resolve these discrepancies, three of the five assessors conducted a second evaluation with only two scoring options for each tied case. This process produced the Labadain-Avaliadór, containing 5,900 qrels, with 9.59% highly relevant documents, 17.86% relevant documents, 9.31% marginally relevant documents, and 63.24% non-relevant documents.
9. Indexing, Retrieval, and Ranking
This section presents the experiments conducted on Tetun ad-hoc text retrieval, providing a detailed description of the retrieval and ranking strategies and the steps involved in text preprocessing, indexing, retrieval, and ranking.
9.1. Overview
The inverted index is one of the most widely used techniques in IR (Baeza-Yates and Ribeiro-Neto, 2011). It is a word-oriented mechanism that indexes all distinct words in the collection, pointing each word to a list of documents in which it appears. This full-text indexing allows direct access to each matching term and its position within the documents.
Studies in ad-hoc text retrieval have demonstrated the effectiveness of various retrieval and ranking models. TF-IDF serves as a foundational term-weighting scheme in IR (Baeza-Yates and Ribeiro-Neto, 2011), and BM25 is a widely recognized probabilistic model known for its effectiveness in classical IR (Robertson and Zaragoza, 2009). Similarly, the probabilistic language model (LM) with Dirichlet smoothing (MacKay and Peto, 1995) has shown strong performance, particularly for short queries in ad-hoc retrieval tasks (Zhai and Lafferty, 2001).
Moreover, the Divergence from Randomness (DFR) variant of BM25 (DFR BM25) has shown competitive performance in various retrieval settings and has been demonstrated in multiple TREC experiments (Wolf et al., 2009; Plachouras et al., 2004). The Hiemstra LM (Hiemstra and Kraaij, 1999; Hiemstra, 2001) has also been reported to perform well in ad-hoc text retrieval, especially in LRL scenarios (Sahu and Pal, 2023). These retrieval and ranking models are used in the experiments conducted in this study.
9.2. Text Preprocessing
Given the language-specific characteristics of Tetun, text preprocessing was divided into several stages as follows:
  1. (1)
    Standard preprocessing: This stage included converting text to lowercase, normalizing apostrophes, removing punctuation and special characters, tokenizing text into tokens (words and numbers), filtering out words longer than 60 characters, and removing extra spaces.
  2. (2)
    Language-specific preprocessing: To address Tetun’s unique linguistic features, additional text preprocessing techniques were applied, including the removal of apostrophes, accents, and hyphens (splitting hyphen-connected words). Each of these steps was independently implemented within the preprocessing workflow.
  3. (3)
    Stopwords Removal and Stemming: Beyond character-based preprocessing, this stage included stopword removal and stemming, with lightmoderate, and heavy variants of the Labadain-Stemmer applied for stemming.
9.3. Experimental Setting
We explored various retrieval strategies by applying multiple text preprocessing techniques to assess their impact on retrieval effectiveness. The experiment workflow is illustrated in Figure 13. First, we established the baseline by applying the standard preprocessing step, as detailed in Subsection 9.2. Next, we tested each of the language-specific text preprocessing techniques, including stopword removal and stemming, to compare their results. Finally, we combined techniques that outperformed the baseline to further evaluate their effectiveness and determine effective strategies for Tetun ad-hoc text retrieval.
Figure 13. Overview of the Experiment Workflow.:
For stemming, although the moderate and heavy variants performed better than the light variant in intrinsic assessment, the difference was minimal. Furthermore, Flores and Moreira (2016) noted that the most accurate stemmer does not always lead to the most effective retrieval, and therefore, we experimented with all stemmer variants. We used PyTerrier (Macdonald and Tonellotto, 2020), a Python API for the Terrier IR platform (Ounis et al., 2005) [^12] http://terrier.org/ for indexing, retrieval, and ranking, with the default settings maintained for each model. The same text preprocessing techniques were applied to both queries and documents, and experiments were conducted using document titles and content to evaluate retrieval effectiveness for each approach.
9.4. Document Indexing
Document titles and content were indexed independently, with the index created from standard preprocessing steps serving as the baseline. Additional indexes were generated for individual results obtained from applying language-specific preprocessing steps, stopword removal, and each of the stemmer variants. To evaluate the effects of these preprocessing methods on index size, index compression factors (ICF) (Frakes and Fox, 2003) were calculated for each preprocessing configuration. A summary of the index compression results is presented in Table 23.
The results show that all preprocessing methods generally reduce the index size compared to the baseline. Removing hyphens yielded the highest compression factor for title indexing, reducing the index size by up to 30.76% compared to the baseline. For content indexing, the heavy stemmer provided the highest compression, reducing the index size by up to 12.18%.
Table 23. Index Compression Factor (%) for Titles and Contents Compared to the Baseline.
Description Baseline No Apostrophes No Accents No Hyphens No Stopwords Light Moderate Heavy
Title 25,412 25,258 0.61 23,568 7.25 17,596 30.76 25,256 0.61 23,416 7.86 23,377 8.01 23,283 8.38
Content 163,203 162,012 0.73 150,596 7.72 146,657 10.14 163,148 0.33 144,240 11.62 143,698 11.95 143,329 12.18
As expected, the moderate stemmer variant compressed the index more efficiently than the light variant, while the heavy variant achieved the highest compression for both title and content indexing. Interestingly, removing apostrophes and accents also contributed to index size reduction, with accent removal achieving over a 7% reduction for both titles and content. This suggests a substantial presence of identical words with and without accents in the documents.
9.5. Short-Text Retrieval Results
For short-text retrieval, document titles were indexed and used for retrieval. The impact of each preprocessing technique on retrieval effectiveness is presented Table 24. Scores highlighted in red indicate values lower than the baseline. As observed, removing accents and applying all stemmer variants did not improve retrieval effectiveness compared to the baseline. Stopword removal yielded inconsistent results across models and metrics. It generally demonstrated improved performance only at the top-20 cutoffs (P@20, MAP@20, and NDCG@20) and overall MAP and NDCG, while exhibiting lower performance at top-5 and top-10 cutoffs across all metrics. Regarding the impact of stemming on retrieval effectiveness, the light stemmer variant slightly outperformed the moderate and heavy variants across all cutoffs, with the moderate and heavy variants performing similarly.
Table 24. Effectiveness of Text Preprocessing Techniques in Short-Text Retrieval. Red values indicate scores lower than the baseline.
Precision at Cutoff MAP at Cutoff NDCG at Cutoff
Retrieval Strategies Model @5 @10 @20 @5 @10 @20 @5 @10 @20 MAP NDCG
BM25 0.8169 0.7763 0.6602 0.1444 0.2568 0.3903 0.6801 0.6668 0.6454 0.5925 0.7408
DFR BM25 0.8169 0.7763 0.6619 0.1440 0.2563 0.3901 0.6811 0.6666 0.6468 0.5926 0.7407
TF-IDF 0.8136 0.7746 0.6458 0.1432 0.2546 0.3825 0.6739 0.6640 0.6380 0.5802 0.7364
Dirichlet LM 0.7898 0.7525 0.6398 0.1299 0.2361 0.3671 0.6359 0.6356 0.6174 0.5780 0.7208
Baseline Hiemstra LM 0.8136 0.7695 0.6669 0.1428 0.2521 0.3928 0.6670 0.6588 0.6465 0.6090 0.7435
BM25 0.8237 0.7763 0.6644 0.1453 0.2572 0.3930 0.6866 0.6685 0.6499 0.5938 0.7419
DFR BM25 0.8237 0.7763 0.6661 0.1450 0.2568 0.3929 0.6878 0.6684 0.6515 0.5942 0.7420
TF-IDF 0.8203 0.7746 0.6500 0.1443 0.2552 0.3854 0.6808 0.6660 0.6428 0.5818 0.7377
Dirichlet LM 0.7898 0.7542 0.6432 0.1301 0.2365 0.3686 0.6380 0.6376 0.6206 0.5794 0.7219
Remove apostrophes Hiemstra LM 0.8169 0.7712 0.6712 0.1429 0.2529 0.3953 0.6725 0.6609 0.6507 0.6102 0.7443
BM25 0.8271 0.7881 0.6856 0.1459 0.2616 0.4069 0.7143 0.7014 0.6871 0.6498 0.8130
DFR BM25 0.8271 0.7881 0.6856 0.1459 0.2616 0.4070 0.7138 0.7016 0.6873 0.6506 0.8135
TF-IDF 0.8271 0.7814 0.6805 0.1457 0.2573 0.4028 0.7118 0.6979 0.6845 0.6402 0.8077
Dirichlet LM 0.7898 0.7576 0.6797 0.1322 0.2420 0.3860 0.6578 0.6615 0.6652 0.6679 0.8039
Remove hyphens Hiemstra LM 0.8339 0.7881 0.6898 0.1472 0.2635 0.4143 0.7142 0.6980 0.6914 0.6841 0.8239
BM25 0.7831 0.7542 0.6424 0.1370 0.2445 0.3735 0.6564 0.6512 0.6308 0.5744 0.7329
DFR BM25 0.7831 0.7542 0.6441 0.1366 0.2440 0.3734 0.6566 0.6509 0.6317 0.5747 0.7328
TF-IDF 0.7864 0.7508 0.6297 0.1364 0.2417 0.3666 0.6555 0.6477 0.6237 0.5631 0.7287
Dirichlet LM 0.7390 0.7237 0.6331 0.1159 0.2173 0.3503 0.5945 0.6081 0.6036 0.5588 0.7104
Remove accents Hiemstra LM 0.7763 0.7441 0.6458 0.1341 0.2388 0.3732 0.6446 0.6406 0.6305 0.5880 0.7352
BM25 0.8102 0.7729 0.6695 0.1438 0.2547 0.3976 0.6693 0.6600 0.6488 0.6030 0.7443
DFR BM25 0.8102 0.7729 0.6712 0.1439 0.2549 0.3984 0.6695 0.6602 0.6503 0.6049 0.7451
TF-IDF 0.8102 0.7729 0.6686 0.1439 0.2549 0.3975 0.6691 0.6600 0.6484 0.6018 0.7438
Dirichlet LM 0.8034 0.7593 0.6653 0.1317 0.2379 0.3803 0.6329 0.6315 0.6299 0.5936 0.7255
Remove stopwords Hiemstra LM 0.8203 0.7678 0.6864 0.1437 0.2521 0.4036 0.6702 0.6587 0.6577 0.6189 0.7483
BM25 0.8000 0.7678 0.6500 0.1381 0.2464 0.3758 0.6693 0.6605 0.6355 0.5826 0.7364
DFR BM25 0.8000 0.7661 0.6492 0.1376 0.2456 0.3748 0.6679 0.6588 0.6345 0.5824 0.7358
TF-IDF 0.7966 0.7610 0.6407 0.1371 0.2421 0.3700 0.6648 0.6551 0.6304 0.5711 0.7331
Dirichlet LM 0.7661 0.7203 0.6305 0.1225 0.2198 0.3518 0.6286 0.6192 0.6076 0.5613 0.7174
Light stemming Hiemstra LM 0.7898 0.7492 0.6576 0.1315 0.2365 0.3760 0.6456 0.6410 0.6326 0.5907 0.7344
BM25 0.7797 0.7610 0.6466 0.1336 0.2423 0.3714 0.6594 0.6553 0.6319 0.5790 0.7346
DFR BM25 0.7797 0.7593 0.6458 0.1331 0.2415 0.3704 0.6581 0.6535 0.6309 0.5789 0.7341
TF-IDF 0.7763 0.7542 0.6373 0.1326 0.2380 0.3656 0.6550 0.6499 0.6268 0.5676 0.7313
Dirichlet LM 0.7593 0.7186 0.6254 0.1197 0.2175 0.3484 0.6225 0.6149 0.6025 0.5577 0.7138
Moderate stemming Hiemstra LM 0.7729 0.7390 0.6525 0.1275 0.2319 0.3710 0.6365 0.6337 0.6277 0.5868 0.7322
BM25 0.7797 0.7610 0.6466 0.1336 0.2423 0.3714 0.6594 0.6553 0.6319 0.5788 0.7346
DFR BM25 0.7797 0.7593 0.6458 0.1333 0.2416 0.3705 0.6583 0.6537 0.6311 0.5788 0.7341
TF-IDF 0.7763 0.7542 0.6373 0.1326 0.2380 0.3656 0.6550 0.6499 0.6268 0.5674 0.7312
Dirichlet LM 0.7559 0.7186 0.6254 0.1189 0.2170 0.3479 0.6200 0.6145 0.6022 0.5571 0.7135
Heavy stemming Hiemstra LM 0.7729 0.7390 0.6534 0.1275 0.2319 0.3715 0.6365 0.6337 0.6282 0.5867 0.7322
Based on the preliminary results in Table 24, all combinations of preprocessing techniques that outperformed the baseline, including stopword removal and the light stemmer variant, were selected for further comparison. The best results from these combined preprocessing techniques are presented in Table 25. The findings indicate that removing apostrophes and hyphens significantly enhances retrieval performance compared to the baseline.
Among the retrieval models, DFR BM25 consistently delivered the highest performance across all metrics up to top-10 cutoffs. Notable scores include P@5 (0.8881), P@10 (0.8390), MAP@5 (0.1589, removing stopwords), MAP@10 (0.2804), NDCG@5 (0.7512), and NDCG@10 (0.7356). DFR BM25 demonstrated relative improvements of up to 30.35% over the baseline and up to 8.19% over individual preprocessing techniques on multiple metrics, along with modest relative gains of up to 5.54% in MAP@5 over other retrieval models within the same settings. While DFR BM25 achieved the same P@5 as BM25 and MAP@5 as TF-IDF, it demonstrated slightly higher scores in other cutoffs.
Table 25. Effectiveness of Combined Text Preprocessing Techniques in Short-Text Retrieval. Values highlighted with a green background indicate the best score at the respective metric and cutoff.
Precision at Cutoff MAP at Cutoff NDCG at Cutoff
Retrieval Strategies Model @5 @10 @20 @5 @10 @20 @5 @10 @20 MAP NDCG
BM25 0.8169 0.7763 0.6602 0.1444 0.2568 0.3903 0.6801 0.6668 0.6454 0.5925 0.7408
DFR BM25 0.8169 0.7763 0.6619 0.1440 0.2563 0.3901 0.6811 0.6666 0.6468 0.5926 0.7407
TF-IDF 0.8136 0.7746 0.6458 0.1432 0.2546 0.3825 0.6739 0.6640 0.6380 0.5802 0.7364
Dirichlet LM 0.7898 0.7525 0.6398 0.1299 0.2361 0.3671 0.6359 0.6356 0.6174 0.5780 0.7208
Baseline Hiemstra LM 0.8136 0.7695 0.6669 0.1428 0.2521 0.3928 0.6670 0.6588 0.6465 0.6090 0.7435
BM25 0.8237 0.7763 0.6644 0.1453 0.2572 0.3930 0.6866 0.6685 0.6499 0.5938 0.7419
DFR BM25 0.8237 0.7763 0.6661 0.1450 0.2568 0.3929 0.6878 0.6684 0.6515 0.5942 0.7420
TF-IDF 0.8203 0.7746 0.6500 0.1443 0.2552 0.3854 0.6808 0.6660 0.6428 0.5818 0.7377
Dirichlet LM 0.7898 0.7542 0.6432 0.1301 0.2365 0.3686 0.6380 0.6376 0.6206 0.5794 0.7219
Remove apostrophes Hiemstra LM 0.8169 0.7712 0.6712 0.1429 0.2529 0.3953 0.6725 0.6609 0.6507 0.6102 0.7443
BM25 0.8271 0.7881 0.6856 0.1459 0.2616 0.4069 0.7143 0.7014 0.6871 0.6498 0.8130
DFR BM25 0.8271 0.7881 0.6856 0.1459 0.2616 0.4070 0.7138 0.7016 0.6873 0.6506 0.8135
TF-IDF 0.8271 0.7814 0.6805 0.1457 0.2573 0.4028 0.7118 0.6979 0.6845 0.6402 0.8077
Dirichlet LM 0.7898 0.7576 0.6797 0.1322 0.2420 0.3860 0.6578 0.6615 0.6652 0.6679 0.8039
Remove hyphens Hiemstra LM 0.8339 0.7881 0.6898 0.1472 0.2635 0.4143 0.7142 0.6980 0.6914 0.6841 0.8239
BM25 0.8102 0.7729 0.6695 0.1438 0.2547 0.3976 0.6693 0.6600 0.6488 0.6030 0.7443
DFR BM25 0.8102 0.7729 0.6712 0.1439 0.2549 0.3984 0.6695 0.6602 0.6503 0.6049 0.7451
TF-IDF 0.8102 0.7729 0.6686 0.1439 0.2549 0.3975 0.6691 0.6600 0.6484 0.6018 0.7438
Dirichlet LM 0.8034 0.7593 0.6653 0.1317 0.2379 0.3803 0.6329 0.6315 0.6299 0.5936 0.7255
Remove stopwords Hiemstra LM 0.8203 0.7678 0.6864 0.1437 0.2521 0.4036 0.6702 0.6587 0.6577 0.6189 0.7483
BM25 0.8881 0.8373 0.7153 0.1553 0.2796 0.4304 0.7500 0.7347 0.7133 0.6648 0.8213
DFR BM25 0.8881 0.8390 0.7169 0.1553 0.2804 0.4313 0.7512 0.7356 0.7149 0.6664 0.8219
TF-IDF 0.8780 0.8322 0.7119 0.1543 0.2759 0.4273 0.7401 0.7288 0.7086 0.6553 0.8149
Dirichlet LM 0.8407 0.8034 0.7068 0.1390 0.2561 0.4099 0.6834 0.6920 0.6829 0.6713 0.8018
Remove apostrophes and hyphens Hiemstra LM 0.8780 0.8305 0.7263 0.1524 0.2743 0.4339 0.7379 0.7245 0.7147 0.6955 0.8282
BM25 0.8814 0.8237 0.7237 0.1576 0.2720 0.4356 0.7394 0.7221 0.7130 0.6752 0.8220
DFR BM25 0.8847 0.8254 0.7237 0.1585 0.2729 0.4366 0.7416 0.7228 0.7139 0.6764 0.8224
TF-IDF 0.8847 0.8237 0.7229 0.1585 0.2722 0.4355 0.7416 0.7220 0.7126 0.6715 0.8202
Dirichlet LM 0.8508 0.8102 0.7220 0.1409 0.2549 0.4171 0.6933 0.6925 0.6918 0.6863 0.8065
Remove hyphens and stopwords Hiemstra LM 0.8746 0.8305 0.7305 0.1524 0.2720 0.4372 0.7294 0.7211 0.7152 0.7040 0.8288
BM25 0.8814 0.8220 0.7212 0.1580 0.2725 0.4342 0.7395 0.7211 0.7123 0.6743 0.8223
DFR BM25 0.8847 0.8237 0.7212 0.1589 0.2735 0.4353 0.7418 0.7218 0.7131 0.6756 0.8228
TF-IDF 0.8847 0.8220 0.7203 0.1589 0.2727 0.4341 0.7417 0.7210 0.7118 0.6705 0.8205
Dirichlet LM 0.8508 0.8068 0.7144 0.1418 0.2546 0.4142 0.6949 0.6913 0.6872 0.6831 0.8053
Remove hyphens, apostrophes, and stopwords Hiemstra LM 0.8746 0.8254 0.7263 0.1528 0.2711 0.4353 0.7287 0.7190 0.7134 0.7029 0.8289
BM25 0.8678 0.8169 0.7110 0.1540 0.2698 0.4267 0.7367 0.7203 0.7053 0.6602 0.8154
DFR BM25 0.8712 0.8153 0.7110 0.1549 0.2695 0.4268 0.7389 0.7200 0.7058 0.6608 0.8157
TF-IDF 0.8712 0.8169 0.7102 0.1549 0.2701 0.4267 0.7389 0.7201 0.7048 0.6584 0.8145
Dirichlet LM 0.8339 0.7932 0.7034 0.1390 0.2503 0.4056 0.6936 0.6860 0.6800 0.6622 0.8023
Remove hyphens and stopwords, and apply light stemmer Hiemstra LM 0.8576 0.8203 0.7229 0.1496 0.2678 0.4289 0.7256 0.7171 0.7083 0.6831 0.8208
Average performance gains of the best model compared to the baseline 8.72% 8.08% 9.54% 10.66% 9.40% 11.30% 10.29% 30.35% 10.63% 15.60% 11.49%
Average performance gains of the best model compared to individual text preprocessing techniques 7.60% 7.39% 6.16% 9.75% 8.19% 6.93% 7.23% 7.45% 6.09% 8.33% 7.58%
Average performance gains of the best model compared to others within the same retrieval strategy 2.65% 1.62% 1.03% 5.54% 3.41% 1.43% 3.35% 2.22% 1.06% 3.94% 1.38%
The removal of hyphens and stopwords also proved beneficial, particularly with Hiemstra LM, which demonstrated the highest scores for P@20 (0.7305), MAP@20 (0.4372), NDCG@20 (0.7152), and overall MAP (0.7040) and NDCG (0.8289, with apostrophe removal). This combination enhanced retrieval effectiveness overall, with Hiemstra LM consistently showing relative improvements at top-20 cutoffs and in global MAP and NDCG. It demonstrated relative improvements of up to 15.60% over the baseline and up to 8.33% over individual text preprocessing techniques in multiple metrics, with marginal relative gains of up to 3.94% over other retrieval models within the same setting.
Stemming, however, had minimal impact on retrieval effectiveness, even when combined with other text preprocessing techniques that outperformed the baseline. The best result involving the light stemmer variant was achieved when paired with hyphen and stopword removal. While this combination produced competitive results, it did not surpass the performance of strategies that excluded stemming.
9.6. Long-Text Retrieval Results
Document content was indexed and used for long-text retrieval. The impact of each preprocessing technique on retrieval effectiveness is summarized in Table 26. The results differ from those of short-text retrieval: apostrophe and accent removal did not provide any measurable benefit for BM25-based models in terms of Precision or MAP relative to the baseline. Likewise, accent removal offered no improvement in P@10, P@20, or MAP with the Dirichlet LM model.
However, there was evidence of improvement with Hiemstra LM in MAP and NDCG at various cutoffs and in overall NDCG. In contrast, hyphen removal, stopword removal, and stemming demonstrated clear advantages for long-text retrieval. Among the stemming variants, the light stemmer consistently outperformed the moderate and heavy variants, though the margin of improvement over the moderate variant was minimal.
Table 26. Effectiveness of Text Preprocessing Techniques in Long-Text Retrieval. Red values indicate scores lower than the baseline.
Precision at Cutoff MAP at Cutoff NDCG at Cutoff
Retrieval Strategies Model @5 @10 @20 @5 @10 @20 @5 @10 @20 MAP NDCG
BM25 0.4847 0.4525 0.3839 0.0769 0.1281 0.1931 0.3883 0.3800 0.3765 0.3429 0.5764
DFR BM25 0.4746 0.4475 0.3839 0.0758 0.1259 0.1925 0.3826 0.3763 0.3753 0.3416 0.5754
TF-IDF 0.5288 0.4746 0.4110 0.0855 0.1382 0.2068 0.4275 0.4086 0.4045 0.3564 0.5927
Dirichlet LM 0.4576 0.4186 0.3669 0.0696 0.1111 0.1701 0.3577 0.3544 0.3545 0.3110 0.5558
Baseline Hiemstra LM 0.5390 0.4915 0.4314 0.0856 0.1402 0.2138 0.4289 0.4158 0.4189 0.3655 0.5990
BM25 0.4881 0.4458 0.3847 0.0775 0.1277 0.1942 0.3902 0.3787 0.3783 0.3441 0.5774
DFR BM25 0.4780 0.4407 0.3839 0.0764 0.1256 0.1934 0.3843 0.3749 0.3765 0.3428 0.5763
TF-IDF 0.5288 0.4695 0.4110 0.0860 0.1388 0.2079 0.4278 0.4088 0.4061 0.3578 0.5938
Dirichlet LM 0.4576 0.4220 0.3661 0.0691 0.1121 0.1707 0.3585 0.3571 0.3549 0.3120 0.5567
Remove apostrophes Hiemstra LM 0.5424 0.4915 0.4297 0.0862 0.1408 0.2149 0.4328 0.4183 0.4200 0.3671 0.6001
BM25 0.5322 0.4966 0.4407 0.0869 0.1437 0.2265 0.4315 0.4270 0.4324 0.4042 0.6570
DFR BM25 0.5288 0.4949 0.4407 0.0860 0.1425 0.2250 0.4259 0.4247 0.4302 0.4026 0.6547
TF-IDF 0.5966 0.5390 0.4686 0.0942 0.1580 0.2425 0.4812 0.4677 0.4649 0.4224 0.6783
Dirichlet LM 0.5051 0.4729 0.4153 0.0813 0.1338 0.2034 0.3961 0.3967 0.4014 0.3729 0.6364
Remove hyphens Hiemstra LM 0.5797 0.5542 0.4907 0.0946 0.1620 0.2503 0.4688 0.4726 0.4771 0.4338 0.6827
BM25 0.4915 0.4441 0.3754 0.0798 0.1283 0.1900 0.3977 0.3815 0.3773 0.3413 0.5798
DFR BM25 0.4780 0.4373 0.3746 0.0784 0.1261 0.1889 0.3899 0.3778 0.3754 0.3396 0.5785
TF-IDF 0.5153 0.4627 0.3966 0.0870 0.1369 0.2013 0.4280 0.4074 0.4011 0.3540 0.5948
Dirichlet LM 0.4542 0.4102 0.3534 0.0657 0.1055 0.1592 0.3517 0.3457 0.3421 0.3037 0.5497
Remove accents Hiemstra LM 0.5424 0.4881 0.4186 0.0861 0.1405 0.2108 0.4382 0.4198 0.4153 0.3634 0.6001
BM25 0.5390 0.4847 0.4212 0.0891 0.1465 0.2182 0.4286 0.4157 0.4156 0.3716 0.6098
DFR BM25 0.5356 0.4864 0.4195 0.0886 0.1463 0.2169 0.4267 0.4155 0.4138 0.3706 0.6088
TF-IDF 0.5559 0.4831 0.4203 0.0924 0.1470 0.2193 0.4423 0.4185 0.4174 0.3716 0.6123
Dirichlet LM 0.4712 0.4271 0.3703 0.0723 0.1176 0.1788 0.3640 0.3553 0.3585 0.3184 0.5656
Remove stopwords Hiemstra LM 0.5458 0.5102 0.4347 0.0895 0.1505 0.2243 0.4341 0.4283 0.4263 0.3796 0.6164
BM25 0.4881 0.4695 0.3992 0.0819 0.1409 0.2096 0.3944 0.3970 0.3949 0.3650 0.6060
DFR BM25 0.4881 0.4712 0.3983 0.0818 0.1410 0.2088 0.3941 0.3979 0.3939 0.3643 0.6056
TF-IDF 0.5525 0.4949 0.4339 0.0925 0.1500 0.2252 0.4449 0.4270 0.4298 0.3810 0.6244
Dirichlet LM 0.4712 0.4441 0.3805 0.0701 0.1197 0.1810 0.3563 0.3616 0.3631 0.3290 0.5762
Heavy stemming Hiemstra LM 0.5458 0.5153 0.4492 0.0868 0.1479 0.2275 0.4312 0.4321 0.4366 0.3871 0.6261
BM25 0.4915 0.4712 0.3992 0.0822 0.1413 0.2099 0.3970 0.3981 0.3953 0.3656 0.6065
DFR BM25 0.4915 0.4712 0.3983 0.0821 0.1411 0.2090 0.3967 0.3982 0.3942 0.3648 0.6060
TF-IDF 0.5525 0.4949 0.4339 0.0925 0.1500 0.2252 0.4449 0.4270 0.4298 0.3814 0.6246
Dirichlet LM 0.4712 0.4441 0.3805 0.0701 0.1197 0.1810 0.3563 0.3616 0.3631 0.3293 0.5764
Moderate stemming Hiemstra LM 0.5458 0.5153 0.4492 0.0868 0.1479 0.2276 0.4312 0.4321 0.4366 0.3874 0.6263
BM25 0.4949 0.4763 0.4008 0.0826 0.1421 0.2106 0.3996 0.4015 0.3963 0.3675 0.6079
DFR BM25 0.4949 0.4763 0.4008 0.0825 0.1420 0.2102 0.3993 0.4016 0.3958 0.3668 0.6075
TF-IDF 0.5559 0.5000 0.4381 0.0929 0.1508 0.2266 0.4475 0.4305 0.4323 0.3837 0.6261
Dirichlet LM 0.4814 0.4508 0.3831 0.0716 0.1216 0.1829 0.3655 0.3701 0.3691 0.3317 0.5817
Light stemming Hiemstra LM 0.5492 0.5203 0.4542 0.0880 0.1496 0.2305 0.4360 0.4369 0.4421 0.3906 0.6301
The results presented in Table 26 were further analyzed by combining text preprocessing techniques to better understand their collective impact on long-text retrieval effectiveness. These combined preprocessing techniques yielded varied impacts across evaluation metrics and models. Hiemstra LM consistently outperformed other models in most combinations and cutoffs, except in P@5 and NDCG@5, where TF-IDF exhibited a slight performance advantage. Notable results for Hiemstra LM include the combination of apostrophe and hyphen removal, which achieved the highest scores at P@10 (0.5576), P@20 (0.4907), MAP@20 (0.2523), and NDCG@20 (0.4790).
Additional configurations, such as incorporating moderate stemming, yielded the best results at MAP@5 (0.1645), while including stopword removal improved NDCG@10 (0.4760) and overall MAP (0.4358). Furthermore, Hiemstra LM showed notable gains in effectiveness when combining hyphen and accent removal, particularly at MAP@10 (0.1645) and NDCG (0.6855). Hyphen removal alone continued to demonstrate its effectiveness in long-text retrieval, achieving the same highest score as its combination with apostrophe removal at P@20 (0.4907).
Table 27. Effectiveness of Combined Text Preprocessing Techniques in Long-Text Retrieval. Values highlighted with a green background indicate the best score at the respective metric and cutoff.
Precision at Cutoff MAP at Cutoff NDCG at Cutoff
Retrieval Strategies Model @5 @10 @20 @5 @10 @20 @5 @10 @20 MAP NDCG
BM25 0.4847 0.4525 0.3839 0.0769 0.1281 0.1931 0.3883 0.3800 0.3765 0.3429 0.5764
DFR BM25 0.4746 0.4475 0.3839 0.0758 0.1259 0.1925 0.3826 0.3763 0.3753 0.3416 0.5754
TF-IDF 0.5288 0.4746 0.4110 0.0855 0.1382 0.2068 0.4275 0.4086 0.4045 0.3564 0.5927
Dirichlet LM 0.4576 0.4186 0.3669 0.0696 0.1111 0.1701 0.3577 0.3544 0.3545 0.3110 0.5558
Baseline Hiemstra LM 0.5390 0.4915 0.4314 0.0856 0.1402 0.2138 0.4289 0.4158 0.4189 0.3655 0.5990
BM25 0.5322 0.4966 0.4407 0.0869 0.1437 0.2265 0.4315 0.4270 0.4324 0.4042 0.6570
DFR BM25 0.5288 0.4949 0.4407 0.0860 0.1425 0.2250 0.4259 0.4247 0.4302 0.4026 0.6547
TF-IDF 0.5966 0.5390 0.4686 0.0942 0.1580 0.2425 0.4812 0.4677 0.4649 0.4224 0.6783
Dirichlet LM 0.5051 0.4729 0.4153 0.0813 0.1338 0.2034 0.3961 0.3967 0.4014 0.3729 0.6364
Remove hyphens Hiemstra LM 0.5797 0.5542 0.4907 0.0946 0.1620 0.2503 0.4688 0.4726 0.4771 0.4338 0.6827
BM25 0.5322 0.4949 0.4407 0.0869 0.1434 0.2271 0.4300 0.4254 0.4328 0.4047 0.6567
DFR BM25 0.5254 0.4932 0.4407 0.0857 0.1423 0.2256 0.4230 0.4232 0.4307 0.4031 0.6543
TF-IDF 0.5966 0.5390 0.4678 0.0944 0.1581 0.2432 0.4788 0.4664 0.4650 0.4232 0.6790
Dirichlet LM 0.5017 0.4729 0.4136 0.0809 0.1348 0.2043 0.3953 0.3983 0.4011 0.3737 0.6369
Remove apostrophes and hyphens Hiemstra LM 0.5831 0.5576 0.4907 0.0955 0.1636 0.2523 0.4721 0.4759 0.4790 0.4355 0.6840
BM25 0.5322 0.4932 0.4280 0.0875 0.1442 0.2222 0.4349 0.4251 0.4273 0.4020 0.6585
DFR BM25 0.5254 0.4932 0.4280 0.0861 0.1439 0.2211 0.4316 0.4272 0.4278 0.4007 0.6578
TF-IDF 0.5831 0.5288 0.4525 0.0948 0.1541 0.2339 0.4744 0.4588 0.4555 0.4175 0.6769
Dirichlet LM 0.5017 0.4576 0.4025 0.0779 0.1261 0.1923 0.3882 0.3847 0.3875 0.3645 0.6288
Remove hyphens and accents Hiemstra LM 0.5831 0.5508 0.4763 0.0955 0.1645 0.2474 0.4812 0.4748 0.4737 0.4330 0.6855
BM25 0.5797 0.5220 0.4619 0.0926 0.1547 0.2361 0.4614 0.4492 0.4521 0.4195 0.6722
DFR BM25 0.5763 0.5203 0.4568 0.0918 0.1549 0.2344 0.4585 0.4483 0.4491 0.4189 0.6718
TF-IDF 0.6102 0.5339 0.4661 0.0961 0.1574 0.2403 0.4849 0.4622 0.4612 0.4232 0.6777
Dirichlet LM 0.4983 0.4695 0.4068 0.0742 0.1276 0.1946 0.3841 0.3860 0.3886 0.3604 0.6235
Remove hyphens and stopwords Hiemstra LM 0.5932 0.5458 0.4797 0.0947 0.1601 0.2451 0.4780 0.4722 0.4714 0.4339 0.6840
BM25 0.5729 0.5153 0.4610 0.0922 0.1540 0.2366 0.4575 0.4457 0.4514 0.4199 0.6726
DFR BM25 0.5695 0.5136 0.4576 0.0913 0.1542 0.2352 0.4545 0.4447 0.4492 0.4193 0.6720
TF-IDF 0.6102 0.5322 0.4661 0.0960 0.1570 0.2410 0.4838 0.4618 0.4608 0.4236 0.6779
Dirichlet LM 0.4881 0.4678 0.4042 0.0731 0.1270 0.1940 0.3786 0.3848 0.3858 0.3611 0.6232
Remove hyphens, apostrophes, and stopwords Hiemstra LM 0.6000 0.5492 0.4797 0.0959 0.1628 0.2471 0.4826 0.4760 0.4727 0.4358 0.6853
BM25 0.5322 0.5017 0.4322 0.0849 0.1459 0.2236 0.4156 0.4176 0.4202 0.3981 0.6450
DFR BM25 0.5254 0.4966 0.4297 0.0844 0.1447 0.2214 0.4110 0.4136 0.4173 0.3968 0.6435
TF-IDF 0.6000 0.5441 0.4636 0.0942 0.1583 0.2364 0.4681 0.4589 0.4561 0.4160 0.6700
Dirichlet LM 0.4915 0.4712 0.4076 0.0788 0.1330 0.1989 0.3795 0.3872 0.3911 0.3663 0.6274
Remove hyphens, apostrophes, and moderate stemmer Hiemstra LM 0.5864 0.5559 0.4805 0.0965 0.1631 0.2475 0.4695 0.4734 0.4702 0.4301 0.6780
Average performance gains of the best model compared to the baseline 15.39% 13.45% 13.75% 12.73% 17.33% 18.01% 13.43% 14.48% 14.35% 19.23% 14.44%
Average performance gains of the best model compared to others within the same retrieval strategy 9.61% 11.77% 11.47% 13.23% 16.40% 12.54% 9.64% 10.15% 11.08% 7.82% 4.65%
In contrast, individual preprocessing techniques, such as accent and apostrophe removal, generally underperformed when applied in isolation, contributing minimally to retrieval effectiveness across different metrics (see Table 26). However, combining these techniques with hyphen and stopword removal, along with moderate stemming, resulted in a strong performance, particularly with Hiemstra LM across most metrics and cutoffs. While the heavy stemming variant slightly outperformed the moderate variant when applied individually, the moderate variant showed a slight advantage when combined with other preprocessing techniques.
Hiemstra LM consistently delivered the highest scores with the combined preprocessing approach across all metrics, except at P@5, where TF-IDF marginally outperformed it. Hiemstra LM demonstrated notable performance improvements over the baseline, achieving gains of up to 13.75% for Precision cutoffs, 18.01% for MAP cutoffs, 14.48% for NDCG cutoffs, 19.23% for overall MAP, and 14.44% for overall NDCG. Additionally, Hiemstra LM achieved an average relative improvement of up to 16.40% over the other retrieval models.
Conversely, BM25 and DFR BM25 did not produce the best results with any preprocessing strategies for long-text retrieval, suggesting that BM25-based models may be less effective in handling long-document contexts. These findings underscore the different impacts of text preprocessing techniques and retrieval models on short- and long-text retrieval tasks in Tetun.
9.7. Discussion
In Tetun ad-hoc text retrieval, text preprocessing techniques that involve the removal of apostrophes and hyphens generally lead to performance gains of up to 8.19% with DFR BM25 in short-text retrieval up to the top-10 cutoffs. When stopword removal is combined with apostrophe and hyphen removal, further gains are observed, particularly at MAP@5 with both DFR BM25 and TF-IDF models and at NDCG with Hiemstra LM. When apostrophes are retained and only hyphens and stopwords are removed, Hiemstra LM shows improvements of up to 8.33% compared to other models at P@20, MAP@20, NDCG@20, and overall MAP. These results suggest that in short-text retrieval, targeting hyphen removal can enhance retrieval effectiveness.
Different strategies for handling stopwords and apostrophes yield varied benefits across retrieval models. Retaining stopwords while removing apostrophes proves more effective with DFR BM25, whereas retaining apostrophes while removing stopwords shows better results with Hiemstra LM at higher cutoffs and overall. Accent removal, however, does not enhance retrieval effectiveness in short-text retrieval, whether applied individually or in combination with other preprocessing techniques.
Removing stopwords does not lead to noticeable improvements in retrieval effectiveness. For instance, at P@20, the relative gain is limited to 0.58% compared to Hiemstra LM with apostrophe and hyphen removal. Similarly, MAP@20 improves by 0.76%, NDCG@20 by 0.07%, and overall MAP by 1.22% relative to the baselines. These findings align with Savoy’s (Savoy, 1999) experiments with French, which showed that retaining stopwords yielded better results with the BM25 model. Similar conclusions were reported by Ghosh and Bhattacharya (2017), who found variability in retrieval effectiveness within the same language and no significant improvement from stopword removal. Additionally, Dolamic and Savoy (2010) observed that retaining stopwords improved retrieval effectiveness in Hindi, with minimal differences for Marathi and Bengali.
Regarding stemming, it does not improve retrieval effectiveness in Tetun, whether applied independently or in combination with other preprocessing techniques. This result is consistent with findings from preliminary experiments on Tetun text retrieval (de Jesus, 2022) and Flores and Moreira’s (Flores and Moreira, 2016) experiments with Dutch, Italian, Spanish, and English. However, studies on other LRLs, including those by Sahu and Pal (2023) (Sanskrit), Sahu et al. (2023) (Urdu), and Adriani et al. (2007) (Indonesian), reported positive impacts of stemming on retrieval effectiveness. This discrepancy likely arises from Tetun’s language-specific characteristics, which may respond differently to stemming than other languages.
While intrinsic evaluations suggested that the moderate and heavy stemmers were slightly more accurate than the light stemmer, these differences did not consistently translate into improved retrieval performance. These findings align with Flores and Moreira’s (Flores and Moreira, 2016) results that higher stemmer accuracy does not always improve retrieval effectiveness.
For long-text retrieval, Hiemstra LM consistently outperformed other models across most metrics and cutoffs. Hyphen removal delivered performance gains across various metrics and models, particularly when combined with apostrophe removal, achieving high scores at P@10, P@20, MAP@20, and NDCG@20. Adding moderate stemming to this combination yielded the highest MAP@5 score, though the improvement was minimal compared to strategies without stemming.
Apostrophe removal showed inconsistent results as an individual preprocessing technique but became more effective when combined with hyphen removal. Similarly, accent removal had varying impacts across models and metrics when applied alone, but combined with other techniques, it achieved a relative improvement of 16.40% in MAP@10 over other models. This outcome aligns with Savoy’s (Savoy, 1999) findings on French, where accent removal slightly improved retrieval effectiveness in long-text retrieval but negatively affected short-text retrieval for several strategies.
9.8. Conclusion
The most effective retrieval strategy for Tetun ad-hoc text retrieval involves short-text retrieval (i.e., using document titles) combined with targeted preprocessing techniques, specifically splitting compound words by removing hyphens and eliminating apostrophes from queries and documents. Together, these techniques enhance both retrieval efficiency and effectiveness. The removal of stopwords shows minimal impact, while preserving accents proves crucial for effective text retrieval in Tetun, likely due to the language’s reliance on diacritics to distinguish word meanings.
Stemming does not improve retrieval effectiveness in short-text retrieval. In long-text retrieval, the moderate stemming variant, when combined with other preprocessing techniques, shows slight improvements at MAP@5; however, the gains remain marginal and are still lower than the best-performing score achieved in short-text retrieval, representing -39.25% relative performance drop. This minimal impact suggests that morphological normalization through stemming is not essential for effective text retrieval in Tetun, likely due to the language’s relatively simple morphology and minimal use of inflectional affixes.
The experimental results show that text preprocessing techniques, particularly the removal of hyphens and apostrophes, are crucial in enhancing the effectiveness of Tetun ad-hoc text retrieval. Among the retrieval models, DFR BM25 performs best for cutoffs up to the top-10 but exhibits greater sensitivity to stopword removal. In contrast, Hiemstra LM provides the most substantial improvements at top-20 cutoffs and beyond, particularly in overall MAP and NDCG.
Overall, this study demonstrates the importance of thoroughly investigating language-specific preprocessing strategies by segmenting them into distinct stages and systematically integrating them to uncover the linguistic components that impact retrieval effectiveness (see Figure 13). It emphasizes the benefit of developing approaches tailored to the unique characteristics of each language, rather than relying solely on established techniques in the literature. The study introduces a detailed and adaptable methodology for establishing ad-hoc text retrieval baselines, particularly for LRLs where such benchmarks have not yet been established. By accounting for the unique morphological and syntactic features of each language, researchers can design more effective and linguistically appropriate retrieval strategies.
10. Conclusions and Future Work
This study presents the development of Tetun text retrieval and establishes the first baselines for the ad-hoc retrieval task. As part of this effort, we created three essential resources: Labadain-Stopwords (de Jesus and Nunes, 2025c), Labadain-Stemmer (de Jesus and Nunes, 2025b), and Labadain-Avaliadór (de Jesus and Nunes, 2025a). These resources are publicly available to the IR and NLP research community under the Creative Commons Attribution-ShareAlike license.
Labadain-Stopwords contains 160 Tetun stopwords, a total comparable to those reported for other low-resource languages (LRLs) in the literature. Labadain-Stemmer is available in three variants—lightmoderate, and heavy—with the light variant specifically targeting Portuguese loanwords commonly used in Tetun. Due to the prominence of these loanwords, the Portuguese stemmer from Snowball was adapted for Tetun by adapting it to the linguistic characteristics of Portuguese-derived words in Tetun. This method can be extended to other LRLs with a similar linguistic nature.
Labadain-Avaliadór consists of 59 topics, 33,550 documents, and 5,900 relevance judgments (qrels), with an average of 36.76 relevant documents per query. This average reflects a balanced representation of the relevant documents in the collection. The balance arises from the fact that the queries were derived from real-world search logs sourced from two platforms: Google Search Console logs for Timor News and logs from searches performed using the Timor News search functionality. This ensures that queries and documents reflect real-world scenarios where the documents exist in the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), which is used as the document collection.
Our investigation involved experimenting with different retrieval strategies, with a focus on the impact of various text preprocessing techniques tailored based on the linguistic characteristics of Tetun. We initially hypothesized that text preprocessing techniques would improve the effectiveness of Tetun text retrieval, and our findings confirmed this hypothesis. In response to our research questions, we conclude that removing hyphens to split compound words into individual words, combined with the removal of apostrophes, enhances retrieval effectiveness overall in both short- and long-text retrieval, with short-text retrieval being the most effective approach for Tetun.
For the retrieval and ranking models, DFR BM25 performs effectively with short-text retrieval up to the top-10 cutoff but shows slightly lower effectiveness when stopwords are removed. Meanwhile, Hiemstra LM consistently demonstrates effective performance across various metrics, particularly beyond the top-10 cutoffs and in overall MAP and NDCG. These findings suggest that Hiemstra LM is more effective for Tetun text retrieval when more than ten documents are prioritized for retrieval. The effectiveness of Hiemstra LM for Tetun text retrieval is consistent with the findings of Sahu and Pal (2023) in their study on Sanskrit, a script-based language spoken in India. Although Tetun uses a Latin script, both languages demonstrate comparable retrieval effectiveness with the Hiemstra LM model in short-text retrieval, measured by MAP. This suggests that Hiemstra LM may adapt effectively to LRLs, regardless of their linguistic or script characteristics.
Future work will investigate semantic search, which captures the contextual meaning of queries and documents rather than relying solely on exact term matches. Integrating large language models (LLMs) into retrieval tasks may open new avenues for enhancing retrieval effectiveness and better aligning retrieval systems with user information needs. Additionally, investigating user search behavior influenced by LLM advancements may reveal evolving patterns in user search intent and information needs, particularly in LRLs like Tetun. Insights from these trends could guide the design of retrieval systems that effectively adapt to changing search behaviors.
11. Acknowledgments
This work is financed by National Funds through the Portuguese funding agency, FCT - Fundação para a Ciência e a Tecnologia, under the Ph.D. studentship grant number SFRH/BD/151437/2021 (DOI: 10.54499/SFRH/BD/151437/2021).
We would like to extend our deepest gratitude to Fanézio Pinto, Altedio Araújo, Roohy Freitas, Alton Freitas, Rita Belo, and Paulo Barreto for their invaluable contributions to the development of Labadain-Stopwords, Labadain-Stemmer, and Labadain-Avaliadór, which serve as key components of this study.
12. Appendices
Subsection 12.1 presents the Labadain-Stopwords, and algorithm of the Labadain-Stemmer is shown in Subsection 12.2.
12.1. The Labadain-Stopwords
The Labadain-Stopwords list (de Jesus and Nunes, 2025c) with their English translations are presented in Table 28.
Table 28. The Labadain-Stopwords.
No. Tetun English No. Tetun English No. Tetun English No. Tetun English
1 an self 41 Hamutuk with, together 81 laek less 121 nunka never
2 aleinde besides 42 hanesan such as 82 lai for a while 122 o you
3 ami we 43 hela remain 83 laiha without 123 oin next/sort/front
4 ami-nia our 44 hikas again 84 lalais quickly 124 oin-oin various
5 antes before, previously 45 hira how much 85 laran inside 125 oinsá how
6 atu so that, to 46 hirak those 86 leten on, in 126 oioin diverse
7 atubele in order to 47 hirak-ne’e these 87 liu exceed, more than 127 oituan few
8 ba to, for 48 ho with 88 liubá ago 128 okos below
9 baibain usuallly, commonly 49 hodi so that 89 liuhosi through 129 ona already
10 bainhira when, while 50 hosi from 90 liuhusi through 130 ou or
11 balu some 51 hotu too, also 91 liuliu especially, particularly 131 para to, in order to
12 barak many 52 hotu-hotu all 92 liután further, more 132 portantu so, therefore
13 bazeia based (on), according (to) 53 husi from 93 loloos exactly, correctly 133 rasik self, own
14 beibeik often, always 54 i and 94 loos very, correct 134 resin over, excess
15 bele be able to, could, may 55 ida a, an 95 lubuk a lot of, many 135 ruma some, any
16 besik near, nearby, almost 56 ida-idak each 96 mai to, toward 136 sai become, out
17 buat thing 57 ida-ne’e this 97 maibé but 137 saida what
18 dala time(s) 58 ida-ne’ebé which one 98 mais however 138 se if, whether
19 dalaruma sometimes 59 iha be, exist 99 maizumenus more or less, approximately 139 who
20 daudauk currently 60 imi you (plural) 100 mak to be 140 sei will, still
21 daudaun currently 61 inklui include 101 maka to be 141 seidauk not yet
22 de’it only, just 62 ita we 102 malu each other 142 sein without
23 depois after that, then, later 63 ita-boot you 103 mas but 143 seluk other
24 dezde since, from 64 ita-nia yours 104 maski despite, although 144 sempre always
25 didi’ak carefully, thoroughly 65 ka or 105 menus less 145 sira they
26 duke than 66 kada each, every 106 mezmu despite, although 146 sira-ne’e these
27 duni indeed 67 karik maybe 107 molok before 147 sira-ne’ebé those who
28 durante during 68 katak that 108 mós also 148 sira-nia their
29 eh or 69 kedas beforehand, immediately 109 nafatin still, remain 149 sira-nian theirs
30 enkuantu while 70 komesa from, begin 110 ne’e this 150 only, unless
31 entaun so, then 71 kona-ba about 111 ne’ebá that 151 tan more
32 entre between, among 72 kotuk behind, last 112 ne’ebé where 152 tanba because
33 entretantu meanwhile 73 kraik below 113 nia he, she 153 tantu so
34 fali again 74 kuandu whenever, while 114 nian of 154 tebes very, so
35 filafali again 75 kuaze almost 115 ninia his, her 155 tenke must
36 foin only just 76 la not 116 ninian his, hers 156 tiha already
37 ha’u I 77 la’ós not 117 no and 157 to’o until
38 ha’u-nia my 78 labele unable, don’t 118 nomós also 158 tomak entire
39 hafoin after 79 ladún not very, not so 119 nu’udar as 159 tuir according to
40 hahú from, begin 80 lae no 120 nune’e like this, in this way 160 uitoan few, a little
12.2. Details of the Labadain-Stemmer Algorithms
The linguistic regions used in the Labadain-Stemmer are shown in Table 29, while Table 30 provides a list of Portuguese-derived suffixes. The algorithm of the Tetun light stemmer variant is detailed in Algorithm 12.2.
Table 29. Linguistic Regions Used in the Labadain-Stemmer.
Region Definition
R1 The region starting after the first non-vowel that follows a vowel, or, if no such non-vowel exists, it is the null region at the end of the word.
R2 The region starting after the first non-vowel that follows a vowel within R1, or, if no such non-vowel exists, it is the null region at the end of the word.
RV If the second letter is a consonant, RV starts after the next vowel. If the first two letters are vowels, it begins after the following consonant. In the case of a consonant-vowel combination, RV starts after the third letter. If none of these conditions are met, RV starts at the end of the word.
Table 30. List of Portuguese-Derived Suffixes.
Suffix Variable Description
eza, ezas, iku, ika, ikus, ikas, izmu, izmus, ável, ível, ista, istas, ozu, oza, ozus, ozas, amentu, amentus, imentu, imentus, adora, adór, asaun, adoras, adores, asoens, ante, antes, ánsia, atória, atóriu, atórias, atórius, amentál general_sufgeneral\_suf A list contains general suffixes
lojia, lojias lojia_suflojia\_suf A list contains loj and lojia suffixes
usaun, usoens usaun_sufusaun\_suf A list contains usaun and usoens suffixes
énsia, énsias ensia_sufensia\_suf A list contains énsia and énsias suffixes
amente amente_sufamente\_suf A string with amente value
iv (appears before the amente suffix) iv_sufiv\_suf A string with iv value
at (takes precedence over the iv, ivaivuivas, or ivus suffixes) at_sufat\_suf A string with at value
oz, ik, ad (presents before the amente suffix) ozikad_sufozikad\_suf A list contains ozik, and ad suffixes
mente mente_sufmente\_suf A string with mente value
ante, avel, ivel (appears before the mente suffix) ante_sufante\_suf A list contains anteavel, and ivel suffixes
idade, idades idade_sufidade\_suf A list contains idade and idades suffixes
abil, is, iv (takes precedence over the idade or idades suffixes) abil_sufabil\_suf A list contains abilis, and iv suffixes
iva, ivu, ivas, ivus iva_sufiva\_suf A list contains ivaivuivas and ivus suffixes
ada, adu, adas, adus, ida, idu, idas, idus, ária, áriu, árias, árius verb_sufverb\_suf A list contains verb suffixes
a, e, i, u, us, as residual_sufresidual\_suf A list contains residual suffixes
Algorithm 1: Tetun Light Stemmer Algorithm
1:$R1$, $R2$, RVRV, word_listword\_list
[…]
17:for all wordword in word_listword\_list do $\triangleright$ Step 1: Word length validation
18:  if $\text{length}(word)<4$ then
19:   Return wordword
[…]
21:   if wordword ends with any suffixsuffix in general_sufgeneral\_suf sorted by length descending then
22:     if Position of suffixsuffix in wordword is in $R2$ then
23:      Delete suffixsuffix from wordword
24:      $stem\leftarrow$ wordword without suffixsuffix
25:      Return stemstem
26:     end if
27:   else if wordword ends with any suffixsuffix in lojia_suflojia\_suf then
28:     if Position of suffixsuffix in wordword is in $R2$ then
29:      Replace suffixsuffix in wordword with lojloj
30:      $stem\leftarrow$ wordword without suffixsuffix concatenates with lojloj
31:      Return stemstem
32:     end if
33:   else if wordword ends with any suffixsuffix in usaun_sufusaun\_suf then
34:     if Position of suffixsuffix in wordword is in $R2$ then
35:      Replace suffixsuffix in wordword with $u$
36:      $stem\leftarrow$ wordword without suffixsuffix concatenates with $u$
37:      Return stemstem
38:     end if
39:   else if wordword ends with any suffixsuffix in ensia_sufensia\_suf then
40:     if Position of suffixsuffix in wordword is in $R2$ then
41:      Replace suffixsuffix in wordword with enteente
42:      $stem\leftarrow$ wordword without suffixsuffix concatenates with enteente
43:      Return stemstem
44:     end if
45:   else if wordword ends with suffixsuffix equals amente_sufamente\_suf then
46:     if Position of suffixsuffix in wordword is in $R1$ then
47:      Delete suffixsuffix in wordword
48:      $stem\_amente\leftarrow$ wordword without suffixsuffix
49:      if stem_amentestem\_amente preceded by iv_sufiv\_suf and the position of iv_sufiv\_suf is in $R2$ then
50:        Delete iv_sufiv\_suf in stem_amentestem\_amente
51:        $stem\_iv\leftarrow$ stem_amentestem\_amente without iv_sufiv\_suf
52:        if stem_ivstem\_iv further proceeded by at_sufat\_suf and the position of at_sufat\_suf is in $R2$ then
53:         Delete at_sufat\_suf in stem_ivstem\_iv
54:         $stem\_at\leftarrow$ stem_ivstem\_iv without at_sufat\_suf
55:         Return stem_atstem\_at
56:        end if
57:        Return stem_ivstem\_iv
58:      else if stem_amentestem\_amente preceded by any suffix_oidsuffix\_oid in ozikad_sufozikad\_suf then
59:        if Position of suffix_oidsuffix\_oid in stem_amentestem\_amente is in $R2$ then
60:         Delete suffix_oidsuffix\_oid in stem_amentestem\_amente
61:         $stem\_ozikad\leftarrow$ stem_amentestem\_amente without suffix_oidsuffix\_oid
62:         Return stem_ozikadstem\_ozikad
63:        end if
64:      end if
65:      Return stem_amentestem\_amente
66:     end if
67:   else if wordword ends with suffixsuffix equals mente_sufmente\_suf then
68:     if Position of suffixsuffix in wordword is in $R2$ then
69:      Delete suffixsuffix in wordword
70:      $stem\_mente\leftarrow$ wordword without suffixsuffix
71:      if stem_mentestem\_mente preceded by any suffix_antsuffix\_ant in ante_sufante\_suf then
72:        if Position of suffix_antsuffix\_ant in stem_mentestem\_mente is in $R2$ then
73:         Delete suffix_antsuffix\_ant in stem_mentestem\_mente
74:         $stem\_ante\leftarrow$ stem_mentestem\_mente without suffix_antsuffix\_ant
75:         Return stem_antestem\_ante
76:        end if
77:      end if
78:      Return stem_mentestem\_mente
79:     end if
80:   else if wordword ends with suffixsuffix equals idade_sufidade\_suf then
81:     if Position of suffixsuffix in wordword is in $R2$ then
82:      Delete suffixsuffix in wordword
83:      $stem\_idade\leftarrow$ wordword without suffixsuffix
84:      if stem_idadestem\_idade preceded by any suffix_ablsuffix\_abl in abil_sufabil\_suf then
85:        if Position of suffix_ablsuffix\_abl in stem_idadestem\_idade is in $R2$ then
86:         Delete suffix_ablsuffix\_abl in stem_idadestem\_idade
87:         $stem\_abil\leftarrow$ stem_idadestem\_idade without suffix_ablsuffix\_abl
88:         Return stem_abilstem\_abil
89:        end if
90:      end if
91:      Return stem_idadestem\_idade
92:     end if
93:   else if wordword ends with suffixsuffix equals iva_sufiva\_suf then
94:     if Position of suffixsuffix in wordword is in $R2$ then
95:      Delete suffixsuffix in wordword
96:      $stem\_iva\leftarrow$ wordword without suffixsuffix
97:      if stem_ivastem\_iva preceded by at_sufat\_suf and the position of at_sufat\_suf is in $R2$ then
98:        Delete at_sufat\_suf in stem_ivastem\_iva
99:        $stem\_at\leftarrow$ stem_ivastem\_iva without at_sufat\_suf
100:        Return stem_atstem\_at
101:      end if
102:      Return stem_ivastem\_iva
[…]
104:   else if wordword ends with suffixsuffix equals verb_sufverb\_suf then
105:     if Position of suffixsuffix in wordword is in RVRV then
106:      Delete suffixsuffix in wordword
107:      $stem\_verb\leftarrow$ wordword without suffixsuffix
108:      Return stem_verbstem\_verb
[…]
110:   else if wordword ends with suffixsuffix equals residual_sufresidual\_suf then
111:     if Position of suffixsuffix in wordword is in RVRV then
112:      Delete suffixsuffix in wordword
113:      $stem\_residual\leftarrow$ wordword without suffixsuffix
114:      Return stem_residualstem\_residual
115:     end if
116:   else
117:     Return wordword
118:   end if
119:  end if
120:end for